### v2

# Unified Triplet-Level Hallucination Evaluation for Large Vision-Language Models

Abstract
Despite the outstanding performance in vision-language reasoning, Large Vision-Language Models (LVLMs) might generate hallucinated contents that do not exist in the given image. Most existing LVLM hallucination benchmarks are constrained to evaluate the object-related hallucinations. However, the potential hallucination on the relations between two objects, i.e., relation hallucination, still lacks investigation. To remedy that, in this paper we design a unified framework to measure object and relation hallucination in LVLMs simultaneously. The core idea of our framework is to conduct hallucination evaluation on (object, relation, object) triplets extracted from LVLMs’ responses, and thus, could be easily generalized to different vision-language tasks. Based on our framework, we further introduce Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark which can be used to study both object and relation hallucination at the same time. We conduct comprehensive evaluations on Tri-HE and observe that the relation hallucination issue is even more serious than object hallucination among existing LVLMs, highlighting a previously neglected problem towards reliable LVLMs. Moreover, based on our findings, we design a simple yet effective training-free approach to mitigate hallucinations for LVLMs, with which, we exceed all open-sourced counterparts on Tri-HE, achieving comparable performance with the powerful GPT-4V. Our dataset and code for the reproduction of our experiments are available publicly at https://github.com/wujunjie1998/Tri-HE.
1 Introduction
Large Vision-Language Models (LVLMs) [1, 2, 3, 4] have attracted significant attention. Despite the superior performances, existing works primarily focus on enhancing the helpfulness of LVLMs without careful consideration of the reliability of responses generated by LVLMs. However, it has already been observed by recent literature that LVLMs suffer from severe hallucination [5, 6, 7, 8, 9], i.e., LVLMs might generate contents that do not exist in the given image, probably due to insufficient training during visual instruction tuning. A typical example is provided in Figure 1(a), where the LLaVA [2] model considers the location to be busy, simply because LLaVA recognizes that it is a train station with several people existing.
With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination. However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination [5, 6, 10] (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning. As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image. Therefore, a unified definition and taxonomy is necessary to integrate different kinds of LVLM hallucination.
2) Hallucination discrimination: existing works [5, 6, 8, 11] rely on self-discrimination (e.g., Yes/No questions) to testify LVLMs’ awareness of objects and relations in the given image. When prompted with an instruction “Is there a {content} in the image?”, hallucination is determinied only when LVLMs answer “Yes.” for a non-existing content. However, these benchmarks require transforming general vision-language tasks into "yes/no" formats, limiting their applicability. Additionally, different LVLMs may have different capability in answering such“yes/no” questions since they are pre-trained on different data, which may make the evaluation results biased [12]. So, here arises our question: can we perform a unified and unbiased evaluation for various kinds of hallucination?
To this end, this paper first proposes a unified framework to measure both object and relation hallucinations (§3). Specifically, it first extracts knowledge graphs as triplets from LVLMs’ responses, then uses external judges to compare these triplets with scene graphs of input images for hallucination evaluation. In this way, our method can provide fine-grained evaluation results on the response to any vison-language tasks. Based on the proposed framework, we then construct Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark to evaluate object and relation hallucination simultaneously (§4). As reported in §5 and Figure 1(b), relation hallucination is indeed a common nightmare even for the powerful GPT-4V model [13], and is even more severe than object hallucination. By comparing the performance of different LVLMs, we further recognize several findings that could potentially decrease LVLM hallucination (§5.2). In addition, our triplet-level hallucination judge empowered by GPT-4 reveals remarkable alignment with human judgments (Table 3). This motivates us to integrate triplet description into LVLMs and propose a simple yet effective training-free hallucination mitigation method (§5.4). With this method, the LLaVA-1.5 model [14] exceeds all the open-sourced counterparts, achieving comparable performance with the powerful GPT-4V, as shown in Figure 1(b) and Table 5.
Our main contributions contain three parts:
  1. 1.
    We introduce a unified framework to measure both the object and relation hallucination of LVLMs simultaneously. Specifically, we propose to conduct a triplet-level hallucination evaluation, which can provide more accurate and fine-grained evaluation results.
  2. 2.
    Based on the framework, we propose Tri-HE, a novel triplet-level fine-grained hallucination evaluation benchmark for LVLMs.
  3. 3.
    We propose a simple yet effective training-free hallucination mitigation method that exceeds all open-sourced counterparts, performing on par with GPT-4V on Tri-HE.
(a) Triplet-level LVLM hallucination evaluation pipeline.
(b) LVLM hallucination comparison.
Figure 1: Overview of Tri-HE evaluation pipeline. (a) With the provision of images, scene graphs, and questions, knowledge graphs (i.e., triplets) are extracted from LVLM responses, which are then judged by GPT-4. (b) The radar plot showcases the evaluation results among different LVLMs (lower values demonstrate fewer hallucinations).
2 Related Work
2.1 Large Vision-Language Models (LVLMs)
The powerful capability exhibited by Large Language Models (LLMs) has facilitated the extension of LLMs towards the multi-modal domain. LLMs are empowered to understand and reason about both images and text by aligning representations from visual encoders to pre-trained language models, followed by visual instruction tuning. LLaVA [14, 2] proposes to use a simple projection layer to integrate the visual representations into textual encoders, which is further enhanced in Shikra [15] by incorporating referential dialogue tasks. Instead, BLIP [16] proposes the Q-Former architecture to extract useful information from the visual representations, which is also used by MiniGPT-4 [3] and InstructBLIP [1]. MoCLE [17] further introduces the Mixture-of-Experts architecture into LVLMs to deal with the data conflict during instruction tuning. Although powerful, existing works primarily focus on improving the helpfulness, without a thorough analysis on the reliability of LVLMs.
2.2 Hallucination Evaluation in LVLMs
With the prevalence of LVLMs, a growing number of studies have been conducted on their hallucination issues [9, 18, 19, 20, 21, 6, 8, 22]. Previous hallucination evaluation works can be categorized into two groups: 1) solely evaluating object hallucinations or do not distinguish different hallucinations [23, 24, 6, 10], which neglects other hallucination types like relation hallucination and is thus not comprehensive. The other type of works use “yes/no” questions to evaluate LVLM’s relation/object hallucinations [5, 25, 8, 11]. However, these benchmarks require transforming general vision-language tasks into “yes/no” formats, limiting their applicability. Also, different LVLMs may have different ability in answering such“yes/no” questions since they are pre-trained on different data, which may bias the evaluation results. To remedy this research gap, our paper proposes a triplet-level evaluation framework that can provide fine-grained object and relation hallucinations for responses to any vision-language tasks, with an evaluation benchmark Tri-HE that incorporates questions requiring more complicated commonsense reasoning.
It is worth noting that a concurrent benchmark, Reefknot [26], also evaluates relation hallucinations at the triplet level. However, compared to Tri-HE, Reefknot has several limitations: 1) It fully relies on existing datasets, which may lead to data contamination, whereas Tri-HE consists of questions synthesized by GPT-4V. 2) Reefknot evaluates only a limited number of relation types, while Tri-HE covers a broader range, enabling more comprehensive evaluation results as shown in §5. 3) Reefknot employs only one entailment-based hallucination discriminator, whereas Tri-HE offers more accurate hallucination detection using GPT-4, which can simultaneously identify both object and relation hallucinations.
3 Tri-HE Framework Formulation
Inspired by the relation extraction [27] tasks in NLP, in this section, we propose a unified framework to evaluate both object and relation hallucinations via the object-relation triplets (i.e., $(\text{Object}_{1},\text{Relation},$ $\text{Object}_{2})$). Here the objects and relations can either be a word or a phrase with attributes. We start by defining object and relation hallucinations via triplets in §3.1, based on which, we define our evaluation metrics and pipeline in §3.2 and §3.3 separately.
3.1 Definitions
As shown in Figure 1(a), given an input image $I$, a question $Q$ related to $I$ with its ground truth answer $A$, and the answer $A_{\theta}$ predicted by a LVLM $A_{\theta}(\cdot|Q,I)$ parameterized with $\theta$, we first define,
  • $G=(V,E)$ as the scene graph of $I$, where $V$ and $E$ refer to all the objects existing in $I$ and all the possible relations among existing objects, respectively.
  • $G^{\prime}=(V^{\prime}\subseteq V,E^{\prime}\subseteq E)$ as the knowledge graph that includes all the required objects and relations to answer $Q$.
  • $G_{\theta}=(V_{\theta},E_{\theta})$ as the knowledge graph extracted from $A_{\theta}$, where $V_{\theta}$ and $E_{\theta}$ include all the objects and all the possible relations among objects mentioned in $A_{\theta}$.
Note that here all graphs can be converted to a set of triplets (i.e., $G=\{(v_{1},e,v_{2})\}$, where $v_{1},v_{2}\in V$ and $e\in E$). A common nightmare in previous LVLM hallucination literature lies in the ambiguous discrimination between prediction hallucinations and errors [28]. To obtain unbiased hallucination evaluation results, we separate them depending on whether or not the wrongly generated objects or relations exist in the given image $I$. Specifically, given a triplet $(v_{1},e,v_{2})\in G_{\theta}$, we have the following definitions,
  • Object hallucination: if $v_{1}\notin V$ or $v_{2}\notin V$, suggesting $A_{\theta}$ includes an object not within $I$. For example, the triplet (location, suggests, popular spot for socializing) in Figure 1(a) encounter an object hallucination since the object “ popular spot for socializing” cannot be obtained from $V$.
  • Relation hallucination: if $v_{1},v_{2}\in V$ yet $e\notin E$, suggesting that $A_{\theta}$ correctly recognizes two related objects from $I$ but pair them with a non-existing relation. For example, the triplet (people, walking around, area) in Figure 1(a) has a relation hallucination since the relation “walking around” cannot be obtained from $G$, despite that the objects are all in $V$.
  • Prediction error: if $v_{1},v_{2}\in V$ and $e\in E$ yet $(v_{1},e,v_{2})\notin G$, suggesting $A_{\theta}$ correctly recognizes objects and relations from $I$, yet pairs in a wrong way.
3.2 Evaluation Metrics
With the above definition in hand, given the knowledge graph $G_{\theta}$ extracted from a model response $A_{\theta}$, we calculate hallucination rates of $A_{\theta}$ as the proportion of hallucinated triplets in $G_{\theta}$. Most previous works (e.g., POPE [5]) directly evaluate the hallucination rate at the object-level with respect to the total number of predicted objects, yet make their results not comparable among LVLMs, since different LVLMs might refer to different numbers of objects in their responses. To address this issue, we instead opt to calculate the hallucination rate in the question- and image-level. Specifically, we calculate two types of hallucination rates, including the question-level hallucination rate ($\text{Hallu}_{\text{Q}}$) and image-level hallucination rate ($\text{Hallu}_{\text{I}}$), as defined in the following,
$\text{Hallu}_{\text{Q}}(\{Q\})=\frac{1}{|\{Q\}|}\left(\sum_{Q^{\prime}\in\{Q\}}\left(\frac{\text{\# HT in}\ G_{\theta}}{\text{\# TT in}\ G_{\theta}}\right)\right)\times 100\%,$ (1)
$\text{Hallu}_{\text{I}}(\{I\})=\frac{1}{|\{I\}|}\left(\sum_{I^{\prime}\in\{I\}}\text{Hallu}_{\text{Q}}(\{Q_{I^{\prime}}\})\right)\times 100\%,$ (2)
where HT is Hallucinated Triplets, TT is Total Triplets, $\{Q\}$ and $\{I\}$ are the sets of questions and images that LVLMs are evaluated on, respectively, and $\{Q_{I^{\prime}}\}\subseteq\{Q\}$ suggest the subsets of questions related to the image $I^{\prime}$. For both metrics, lower values demonstrate fewer hallucinations. Since the total number of questions and images is maintained the same for all evaluated LVLMs, $\text{Hallu}_{\text{Q}}(\cdot)$ and $\text{Hallu}_{\text{I}}(\cdot)$ are indeed comparable and unbiased.
3.3 Evaluation Pipeline
With the definitions and evaluation metrics provided in §3.1 and §3.2, the remaining problems contain two parts: 1) how to extract the knowledge graph $G_{\theta}$ from LVLM responses $A_{\theta}$, and 2) how to judge a triplet in $G_{\theta}$ is hallucinated or not. The overview of our pipeline is illustrated in Figure 1(a).
Knowledge Graph Extraction.
Given an LVLM response $A_{\theta}$ with the corresponding question $Q$ and image $I$, we extract the knowledge graph $G_{\theta}$ from $A_{\theta}$ via prompting GPT-4. Check our prompt for knowledge graph extraction in Figure 6. Afterward, we propose two different strategies to judge whether a triplet $(v_{1},e,v_{2})\in G_{\theta}$ includes hallucination based on the ground truth answer $A$ and the image scene graph $G$, as described in the following.
Figure 2: Visualization of data samples in Tri-HE. Each image is associated with a scene graph and question-answer pairs with the reasoning triplet annotations.
NLI Judge.
The first strategy is implemented with a natural language inference (NLI) [29] model  [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Specifically, given an extracted triplet, we first calculate its cosine similarity scores with all triplets in the image scene graph $G$ and only retain those ground truth (GT) triplets with similarity scores greater than 0.5 to refine the information that will be used for the NLI model. If no triplets in $G$ meet this criterion, only the top three GT triplets with the highest similarity scores will be kept, which are then taken as ground truth inputs for the NLI model to make predictions. If the NLI score between the extracted triplet and ground truth triplets is lower than 0.6, suggesting the extracted triplet cannot be induced based on GT triplets, and therefore, resulting in a hallucination.
# Images # Questions # Objects # Relations # Questions / Image # Triplets / Scene Graph
300 1226 1723 618 4.09 19.10
Table 1: Statistics of Tri-HE.
GPT-4 Judge.
The other strategy is via prompting the GPT-4 model, which is a common practice for recent LLM evaluation literature [30]. We prompt GPT-4 to justify whether a previously extracted triplet $(v_{1},e,v_{2})\in G_{\theta}$ can be directly obtained or inferred based on the image scene graph $G$. Note that we do not adopt the GPT-4V judge, since [31] have observed that the text-only GPT-4 judge better aligns with the human preference than GPT-4V. Moreover, if $(v_{1},e,v_{2})$ is recognized as a hallucinated triplet, we further require it to tell whether the hallucination relates to the relation $e$ or objects $v_{1},v_{2}$. Check our prompt for GPT-4 judge in Figure 8  [^2] For both knowledge graph extraction and GPT-4 judge, we use the “gpt-4-1106-preview” model via OpenAI’s API calls with its default inference parameters..
With triplet-level hallucination judgments in hand, we can adopt Eq. (1) and Eq. (2) to calculate the corresponding hallucination rates of $A_{\theta}$, which can be further separated into object hallucination and relation hallucination rates if GPT-4 judge is adopted. Check more detailed results in Table 2.
4 Tri-HE Construction
Following the formulation in §3, in this section, we provide a detailed discussion on how to construct our Tri-HE benchmark for a unified triplet-level evaluation of both hallucinations in LVLMs.
Dataset.
We build Tri-HE upon the GQA dataset [32], which requires LVLMs to conduct commonsense reasoning based on a given image. The provision of scene graph annotations aligns with our formulation of hallucination evaluation at the triplet-level.
Data collection.
However, several scene graphs in GQA are only equipped with parts of the relations among objects appearing in the image, making necessary information required for question answering absent. To address that, we first employ a filtering process to identify images with adequate distinct relations in the scene graph annotation. Specifically, we start by selecting 300 images with more than 5 relations from the corresponding GQA scene graph annotations. To enlarge the question-answer pairs for each GQA image, we utilize GPT-4V  [^3] We use the “gpt-4-vision-preview” model here, as in §5.2. to generate new questions requiring commonsense reasoning based on the input image. By leveraging the prompt provided in Figure 7, we prompt GPT-4V to generate ten questions for each image along with their corresponding answers and relation triplets that illustrate the reasoning process. Afterward, we manually verify the correctness of generated questions, answers, and triplets, while only questions indeed requiring reasoning are maintained. Moreover, the verified triplets are added to the original scene graphs collected in GQA.
Method GPT-4 Judge NLI Judge
Overall Object Relation Overall
$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
MiniGPT-4 53.60 51.79 28.32 26.77 25.25 24.98 55.61 53.36
InstructBLIP 46.68 45.57 22.19 20.88 24.50 24.69 58.25 55.56
LLaVA 42.34 41.30 19.88 18.50 22.46 22.80 54.49 51.51
Shikra 42.20 41.76 18.55 17.54 23.65 24.22 56.46 53.98
LLaVA-1.5 40.66 39.10 18.63 17.28 22.03 21.82 54.14 51.67
LLaMA-3.2 40.16 38.95 22.30 21.08 17.86 17.87 48.46 45.64
InternLM2 38.83 37.54 18.25 17.50 20.58 20.04 54.41 52.08
Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM2 is short for InternLM-XComposer2 [33]
Statistics.
The overall statistics for Tri-HE are summarized in Table 1. As described in Figure 2, each image in Tri-HE is linked to a scene graph and a set of question-answer pairs that require reasoning, accompanied by ground truth triplet annotations. Note that since the quality of each question in Tri-HE is manually verified, expanding its size requires significant resources and poses challenges. Nonetheless, the number of images and questions in Tri-HE is comparable to existing LVLM hallucination evaluation benchmarks such as [23] and [8]. Furthermore, as demonstrated in §5, Tri-HE is able to produce reliable hallucination evaluation results.
5 Results on the Tri-HE Benchmark
5.1 Evaluated LVLMs
We selected six open-source LVLMs for evaluation, including the LLaVA series [14, 2], MiniGPT-4 [3], InstructBLIP [1], Shikra [15], and InternLM-XComposer2 (abbrev., InternLM2) [33]. For all evaluated LVLMs, we selected the 7B variants to ensure fair comparison. Additionally, we test the recent popular Llama-3.2-Vision-Instruct model (abbrev., LLaMA-3.2) [34] and used its smallest version (11B). The official prompt templates and inference configurations were used for all LVLMs. All experiments are conducted on two Nvidia A100 GPUs.
5.2 Main Result
LVLM comparison.
Table 2 compares hallucination rates of different LVLMs on our Tri-HE benchmark. As can be seen, all the evaluated LVLMs suffer from generating hallucinations with at least 38% hallucination rates. Among these LVLMs, InternLM2 [33] obtains the best overall performances, suggesting that its strategy to train with both text-image and textual-only instruction data simultaneously helps better align its visual encoder and LLM, and thus, reduces its hallucination rates. Moreover, compared to LLaVA [2], Shikra [15] has consistently lower hallucination rates, which is built upon LLaVA’s structure with extra grounding capability introduced, indicating that introducing extra grounding could help LVLMs reduce hallucination. Additionally, LLaMA-3.2 achieves the lowest relation hallucination rates, suggesting that a strong textual backbone can help mitigate relation hallucination. However, it exhibits a weaker ability to accurately identify objects, impacting its object and overall hallucination rates. Since LLaMA-3.2 does not outperform other LVLMs with even more parameters, we do not adopt it in the remaining.
Relation hallucination is more severe.
Except for MiniGPT-4 and LLaMA-3.2, all the LVLMs generate more relation hallucinations than object hallucinations. A possible explanation is that existing LVLMs lack reasoning abilities, which makes them easily confused and mess up the relations among objects. This further suggests that focusing on object hallucination [5] is not enough for a throughout analysis of the LVLM reliability and a unified and comprehensive study like our proposed triplet-level evaluation is necessary.
Evaluation pipeline.
In addition, we observe that GPT-4 judge can provide clearer and more reasonable discrimination between models compared to NLI judge. We provide a more comprehensive investigation into the differences between these two judges later in §5.3. Besides, the evaluation results under both $\text{Hallu}_{\text{I}}$ and $\text{Hallu}_{\text{Q}}$ metrics demonstrate the same trend, proving the robustness of our proposed triplet-level hallucination evaluation setting under different evaluation granularities.
Method LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM2 GPT-4V
NLI Judge (Sentence) 0.2182 0.0970 0.3609 0.2596 0.2684 0.2524 0.2787
NLI Judge (Triplet) 0.2951 0.2838 0.2264 0.4259 0.2829 0.2647 0.4190
GPT-4 Judge (Sentence) 0.6631 0.5409 0.3669 0.5532 0.5821 0.5998 0.5548
GPT-4 Judge (Triplet) 0.8115 0.6320 0.4283 0.6235 0.6939 0.7169 0.7292
Table 3: Pearson correlation scores among automatic hallucination judgments and human judgments. The best results under each column are boldfaced.
LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM2
Original $\text{Hallu}_{\text{I}}\downarrow$ 22.46 22.03 25.25 24.50 23.65 20.58
$\text{Hallu}_{\text{Q}}\downarrow$ 22.80 21.82 24.98 24.69 24.22 20.04
First 20% $\text{Hallu}_{\text{I}}\downarrow$ 20.86 18.44 23.00 21.73 22.47 18.57
$\text{Hallu}_{\text{Q}}\downarrow$ 18.73 18.06 22.68 19.82 19.34 16.10
Table 4: Relation hallucination rates for the top 20% frequent object pairs of different LVLMs under the GPT-4 judge. Original refers to the results in Table 2.
Performance of GPT-4V.
Except for the open-sourced LVLMs, we also explore the performance of state-of-the-art commercial LVLMs [13, 35]. Due to the experiment budget, here we only evaluate the GPT-4V [13] model with 25 randomly selected samples in Tri-HE. Specifically, we directly prompting GPT-4V to gather its answers and calculate the corresponding hallucination rates, following the same setting in Table 2. We also report the performances of open-sourced LVLMs on the 25 selected images. As can be seen in Figure 1(b), GPT-4V demonstrates its superior capability and outperforms all open-sourced LVLMs. Although its object hallucination rates are slightly higher than InternLM2 since it tends to associate many objects that are not in the image, GPT-4V has significantly lower reasoning hallucination rates than other LVLMs, indicating its strong reasoning ability, and therefore, reports superior overall hallucination rates.
5.3 Analysis
Investigating automatic hallucination judgments with human judgments.
In §3, we propose to measure hallucination on triplet-level and design two automatic hallucination judges. Here, we further illustrate the effectiveness of the triplet-level evaluation setting by studying its correlation with human judgments. To conduct fine-grained hallucination analysis, previous works [36, 37] split a model response into sub-sentences first, on which their hallucination measurements are conducted. We regard this method as a baseline for comparison. Specifically, we sample a subset of 20 images from Tri-HE and invite human annotators to score five-point-scale hallucination rates of the responses of all the LVLMs in §5.1. Check Table 9 for the detailed annotation guidelines. The human annotators achieve a Krippendorff’s alpha score [38] of 0.66, indicating a high inter-agreement.
Results are shown in Table 3. We find that triplet-level hallucination rates have higher correlations with human judgments with both NLI and GPT-4 judges, indicating that identifying hallucination on triplets can lead to a more accurate, human-preferred evaluation for model responses. Moreover, we notice that the GPT-4 judge achieves a higher correlation to human judgments compared to the NLI counterpart, revealing GPT-4’s superior ability to find hallucinations, which is also consistent with our observation in §5.2.
Figure 3: An illustration demonstrating hallucination mitigation. The three prompting strategies (w/o Mitigation, General Description, and Triplet Description) are listed from left to right. Hallucinated content is highlighted in Red and repeating contents are marked with italic and underline.
Investigating relation hallucination with object information.
As concluded from §5.2, existing LVLMs tend to generate both object and relation hallucinations in their replies, while the relation hallucination rates are even higher. Since different LVLMs have pairs of objects $(v_{1},v_{2})$ that they are familiar with (e.g., high-frequency object pairs in the instruction data they are fine-tuned on) and might generate correct relations on these objects easily, we suppose that the relation hallucination problem might mostly be located in less-frequent object pairs. To verify this assumption, we extract all object pairs for each LVLM from their respective $G_{\theta}$ generated from responses on Tri-HE, and rank these pairs based on their frequency  [^4] Details of this process can be seen in Appendix D.. Then, we calculate each LVLM’s relation hallucination rates on their most frequent object pairs.
As in Table 4, all the LVLMs have significantly lower relation hallucination rates on frequent object pairs they are familiar with, suggesting that they know the possible relations among objects and understand how to choose a relation appropriately. Check more analyses regarding the impact of input length on hallucination rates in Appendix A.
Mitigation GPT-4 Judge NLI Judge
$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
w/o Mitigation 45.86 47.44 55.93 54.94
General Description 46.50 49.19 54.59 53.03
Triplet Description 44.14 42.96 51.19 47.12
Table 5: Hallucination mitigation results on MiniGPT-4. The best results under each column are boldfaced.
5.4 Hallucination Mitigation
Regarding the reasons of LVLM hallucination, previous studies [36, 39] have identified modality misalignment as a key factor. Inspired by ECSO [40], we propose a training-free LVLM hallucination mitigation method via self-alignment [12, 40, 41]. Specifically, we disable the direct visual access of LVLMs to alleviate hallucinations resulting from modality misalignment and fully leverage instruction-following capabilities of LVLMs’ (text-only) LLM backbone for hallucinations mitigation [24].
Method.
We start by prompting LVLMs to describe the image on awareness of the given question (i.e., General Description). Subsequently, we prevent LVLMs from having direct access to the image to downgrade the text-image LVLMs to text-only LLMs, which are then encouraged to respond only based on the previously generated image description but without access to the image.
As discussed in §5.3, better hallucination evaluation is observed at the triplet level. Thus, we further prompt LVLMs to focus more on identifying objects and inter-object relations (i.e., Triplet Description) as triplets when describing the image. We use the subset in Figure 1(b) with MiniGPT-4, with the prompts and an example in Figure 3.
Results.
As shown in Table 5, our results demonstrate the effectiveness of hallucination mitigation through our proposed prompting strategies. Both general and triplet descriptions achieve a significant hallucination rate reduction while prompting LVLMs to focus on objects and inter-object relations achieves the lowest hallucination rates. Additionally, we conduct ablation studies and compare our method with existing baselines in Appendix B.
6 Conclusion
Starting from a unified definition of hallucinations, we propose a novel triplet-level LVLM hallucination evaluation framework for both object and relation hallucinations. Then we introduce Tri-HE, a novel triplet-level LVLM hallucination evaluation benchmark, with which, we conduct a throughout analysis of the discrepancy among object and relation hallucinations. Finally, we propose a simple yet effective training-free hallucination mitigation method, which integrates our findings regarding objects and inter-object relations.
Limitations
Due to the experiment budget, in Figure 1(b), we evaluate GPT-4V on a randomly selected subset of 25 images; in Table 3, we conduct human annotation on another randomly selected subset of 20 images. Although we do not interfere with the data selection process, there may still exist a few biases in the corresponding results. Also, we do not have the budget to evaluate more LVLMs on Tri-HE. We will try to provide more comprehensive results in future works once we get more experimental budgets.
Appendix
Appendix A More Analyses
Investigating hallucination rates with response length.
Previous studies on LVLM hallucination evaluation suggest that the length of model responses may influence the extent of hallucination [5, 39], as some LVLMs tend to produce shorter, safer outputs. However, directly instructing an LVLM to generate a response of a specific length is challenging. To address this, we instead truncate the responses to the first $K$ tokens and compute hallucination rates, varying $K$ to assess its impact on the results.
As shown in Figure 4, while the exact hallucination rates vary, the ranking of different LVLMs remains consistent as the number of tokens increases from 10. Overall, as fewer tokens provide insufficient data for triplet extraction, this finding supports the robustness of our proposed triplet-level evaluation across LVLMs with varying response lengths.
Figure 4: Trends of the hallucination rates of the image-level (left) and question-level (right) evaluations for different LVLMs with respect to the number of tokens in the model responses.
Appendix B More on Hallucination Mitigation
Ablation study.
We conducted an ablation study on a general description (i.e., “describe the image”). As shown in Table 6, the combination of both triplet description prompting and disability of visual input contributed to the effectiveness of our hallucination mitigation approach.
Mitigation GPT-4 Judge NLI Judge
$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
w/o Mitigation 45.86 47.44 55.93 54.94
Ours (eyes-close) 46.50 49.19 54.59 53.03
Ours (triplet) 45.65 45.16 59.35 55.57
Ours (triplet+ eyes-close) 44.14 42.96 51.19 47.12
Table 6: Ablation study on MiniGPT-4 [3]. The best results under each column are boldfaced.
Mitigation on LLaVA-1.5.
To provide a comprehensive evaluation, we conducted additional experiments using another LVLM (i.e., LLAVA-1.5) to mitigate hallucinations. The results presented in Table 7 indicate that both the general description and triplet description methods show an improved reduction in hallucinations, further highlighting the effectiveness of our mitigation approach.
Mitigation GPT-4 Judge NLI Judge
$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
w/o Mitigation 30.72 30.17 53.84 52.06
General Description 28.70 29.80 51.40 49.80
Triplet Description 28.39 32.68 48.97 48.40
Table 7: Further experiment conduct on LLaVA-1.5 [14]. The best results under each column are boldfaced.
Baseline Comparison.
Current methods for mitigating hallucinations typically involve retraining, integrating external detection models, and devising decoding strategies. Compared to existing works, our approach is a plug-and-play method that neither requires costly retraining nor relies on external models. To make a more fair comparison, we conducted experiments with LogicCheckGPT [42], a training-free approach that addresses hallucinations through prompting interactions with the help of GPT-3.5. Under the cost consideration, an evaluation was conducted only with the NLI judge. The results indicate that our method outperforms LogicCheckGPT, highlighting its effectiveness in mitigating hallucinations as shown in Table 8.
Model Mitigation $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
LLaVA-1.5 w/o Mitigation 53.84 52.06
LogicCheckGPT 51.10 50.84
Ours (triplet+ eyes-close) 48.97 48.40
MiniGPT-4 w/o Mitigation 55.93 59.94
LogicCheckGPT 52.34 53.04
Ours (triplet+ eyes-close) 51.19 47.12
Table 8: Hallucination mitigation results on LLaVA-1.5 [14] and MiniGPT-4 [3] with baseline comparison. The best results under each column are boldfaced.
Qualitative comparison
with LogicCheckGPT [42] is provided in Figure 5.
Figure 5: More illustrations demonstrating hallucination mitigation with comparison to LogicCheckGPT. Hallucinated content is highlighted in Red.
Appendix C NLI Threshold Selection
We randomly selected question instances from 10 images and reviewed the set of filtered triplets that were returned. The similarity score threshold was adjusted to 0.5 for the most reasonable returned triplets. These triplets later concatenate together as the ground truth required for generating NLI judgments. In determining if a generated triplet was hallucinated, we further review the NLI judgment results in different thresholds, ultimately deciding on a threshold of 0.6.
Appendix D Object Pairs Extraction and Ranking
In this section, we detailedly describe how we obtain object pairs and their rankings from LVLM responses. Suppose we have all an LVLM’s responses to all questions in Tri-HE, i.e., $G_{\theta}$, we first extract all the object pairs $(v_{1},v_{2})$ from $G_{\theta}$. Then for each object, we replace it by the name of its synset using WordNet to reduce the total types of objects. Afterward, we could calculate the frequency of each object pair and rank them based on their frequency. This ranking will then be used to calculate the first 20% frequent object pairs in Table 4.
Appendix E More on Prompts
E.1 Prompt for triplets extraction with GPT-4
The prompt for extracting triplets in the answer generated by LVLMs is illustrated in Figure 6.
Given a description of the image, please extract a KG from the text and represent the KG with triples formatted with ("subject", "predicate", "object") with readability, each triplet in a line. If ’and’ or ’or’ exists in the input sentence, split the objects into multiple triplets. Please do not extract from uninformative sentences. Here are some in-context examples: ### Input: Optimus (or Tesla Bot) is a robotic humanoid under development by Tesla, Inc. It was announced at the company’s Artificial Intelligence (AI) Day event on August 19, 2021. It is planned to measure 5 ft 8 in (173 cm) tall and weigh 125 lb (57 kg). It is hard to answer whether Tesla is good to invest without more information. ### KG: ("Optimus", "is", "robotic humanoid") ("Optimus", "under development by", "Tesla, Inc.") ("Optimus", "also known as", "Tesla Bot") ("Tesla, Inc.", "announced", "Optimus") ("Announcement of Optimus", "occured at", "Artificial Intelligence (AI) Day event") ("Artificial Intelligence (AI) Day event", "held on", "August 19, 2021") ("Artificial Intelligence (AI) Day event", "organized by", "Tesla, Inc.") ("Optimus", "planned to measure", "5 ft 8 in (173 cm) tall") ("Optimus", "planned to measure", "weigh 125 lb (57 kg).") <Done> ### Input: The image doesn’t provide information about the popularity of the song. The song "Here Comes the Boom" was originally released by American rock band Nelly in 2002 for the soundtrack of the film "The Longest Yard." KG: ("The song ’Here Comes the Boom’", "originally released by", "American rock band Nelly") ("The song ’Here Comes the Boom’", "released in", "2002") ("The song ’Here Comes the Boom’", "featured in", "soundtrack of the film ’The Longest Yard’") ("American rock band Nelly", "released", "The song ’Here Comes the Boom’") ("The Longest Yard", "had soundtrack featuring", "The song ’Here Comes the Boom’") <Done> Now generate the KG for the provided input text: ### Input: {input_text} ### KG:
Figure 6: Prompt for triplets extraction with GPT-4.
E.2 Prompt for question generation with GPT-4V
The prompt for generating questions, answers, and corresponding triplets with GPT-4V is shown in Figure 7.
Generate ten questions about the given image that require an inferential answer, which is not directly observable from the image. The answer to each question can be explained by one or more (object, relation, object) triplets that appear in the scene graph of the given image. Note that the triplets should consist of objects and relations that are visible in the given image. Output the results in the format of: Generated Questions: Answers: Explanations:
Figure 7: Prompt for question generation with GPT-4V.
E.3 Prompt for GPT-4 Judge
The prompt for our proposed GPT-4 judge method is illustrated in Figure 8.
Given a list of reference triplets ("object1", "relation", "object2") extracted from the scene graph of an image, along with a list of objects observed in this image, your task is: Task 1. Determine if a claim triplet ("object1", "relation", "object2") is directly supported by any single triplet in the reference, or can be logically inferred from multiple reference triplets and the list of objects. Follow these steps when finishing the task: 1. Answer "yes" if the claim appears in the reference. 2. Answer "yes" if the claim can be logically inferred from one or more triplets in the reference. Consider: a. General Inferences: Assess common associations or implications. b. Conditional Phrases: Note phrases like "could be", "might", "suggests", which allow broader inferences. c. Equivalence of Objects: In your judgment, treat objects of the same kind as equal. For example, "woman", "man" should be considered under the general category of "person". d. Support from Object List: If the claim is not directly supported or inferable from the triplets, assess whether the list of objects provides additional evidence to support or infer the claim. 3. Answer "no" if the claim neither directly matches any triplet in the reference nor can be reasonably inferred from the triplets and the object list. Task 2: Error categorization. If your answer to the previous task is "no", determine whether the not supported/inferred part in the claim is "object1" or "object2" or "relation". Reference: <REFERENCE> List of Objects: <LIST_OF_OBJECTS> Claim: <CLAIM>
Figure 8: Prompt for the GPT-4 Judge method.
E.4 Prompt for LVLMs in Evaluation
The prompt for generating responses from LVLMs for evaluation is the combination of a question and the corresponding image.
Appendix F Human Annotation Guideline
The detailed guidelines of our human evaluation tasks are shown in Table 9. Noting that two types of inferences in the model responses are regarded as hallucinations during human annotation:
  1. 1.
    Unreasonable inferences (inferences that violate commonsense knowledge).
  2. 2.
    Inferences that are correct, yet cannot be correctly inferred from the image.
Score Description
1 1) The text is totally hallucinated, and is irrelevant to the given input image and question.
or
2) The text is very hard to understand.
2 1) Most of the given responses are hallucinated, yet few sentences of them (one or two) are related to the given image and question.
3 1) Half of the sentences in the given response are hallucinated.
4 1) Most of the sentences in the generated response are not hallucinated.
5 1) No hallucination exists in the generated response.
Table 9: Detailed human evaluation instructions.
Appendix G More Discussion
Future works.
Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models [43, 44, 45, 46, 47, 48] is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder [49, 50, 51, 52] and visual tools (e.g., object detectors [53, 54]) to better extract visual information for LLM reasoning.

### v3

@@
 However, the potential hallucination on the relations between two objects, i.e., relation hallucination, still lacks investigation.
-To remedy that, in this paper we design a unified framework to measure object and relation hallucination in LVLMs simultaneously.
-The core idea of our framework is to conduct hallucination evaluation on (object, relation, object) triplets extracted from LVLMs’ responses, and thus, could be easily generalized to different vision-language tasks.
+To remedy that, we design a unified framework to measure the object and relation hallucination in LVLMs simultaneously.
+The core idea of our framework is to evaluate hallucinations via (object, relation, object) triplets extracted from LVLMs’ responses, making it easily generalizable to different vision-language tasks.
 Based on our framework, we further introduce Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark which can be used to study both object and relation hallucination at the same time.
-We conduct comprehensive evaluations on Tri-HE and observe that the relation hallucination issue is even more serious than object hallucination among existing LVLMs, highlighting a previously neglected problem towards reliable LVLMs.
-Moreover, based on our findings, we design a simple yet effective training-free approach to mitigate hallucinations for LVLMs, with which, we exceed all open-sourced counterparts on Tri-HE, achieving comparable performance with the powerful GPT-4V.
+With comprehensive evaluations on Tri-HE, we observe that the relation hallucination issue is even more serious than object hallucination among existing LVLMs, highlighting a previously neglected problem towards reliable LVLMs.
+Moreover, based on our findings, we design a simple training-free approach that effectively mitigates hallucinations for LVLMs.
 Our dataset and code for the reproduction of our experiments are available publicly at https://github.com/wujunjie1998/Tri-HE.
 1 Introduction
-Large Vision-Language Models (LVLMs) [1, 2, 3, 4] have attracted significant attention.
+Large Vision-Language Models (LVLMs) (Dai et al., 2023; Liu et al., 2023b; Chen et al., 2024a; Cai et al., 2024) have attracted significant attention.
 Despite the superior performances, existing works primarily focus on enhancing the helpfulness of LVLMs without careful consideration of the reliability of responses generated by LVLMs.
-However, it has already been observed by recent literature that LVLMs suffer from severe hallucination [5, 6, 7, 8, 9], i.e., LVLMs might generate contents that do not exist in the given image, probably due to insufficient training during visual instruction tuning.
-A typical example is provided in Figure 1(a), where the LLaVA [2] model considers the location to be busy, simply because LLaVA recognizes that it is a train station with several people existing.
+However, it has already been observed by recent literature that LVLMs suffer from severe hallucination (Li et al., 2023e; Wang et al., 2023b; c; Guan et al., 2024; Chen et al., 2024b), i.e., LVLMs might generate contents that do not exist in the given image, probably due to insufficient training during visual instruction tuning.
+A typical example is provided in Figure 1(a), where the LLaVA (Liu et al., 2023b) model considers the location to be busy, simply because LLaVA recognizes that it is a train station with several people existing but without reasoning about their relationships.
+(a) Triplet-level LVLM hallucination evaluation pipeline.
+(b) LVLM hallucination comparison.
+Figure 1: Overview of the unified hallucination evaluation pipeline of Tri-HE.
+(a) With the provision of images, scene graphs, and questions, knowledge graphs (i.e., triplets) are extracted from LVLM responses, which are then judged by an LLM (GPT-4 here).
+(b) The radar plot showcases the evaluation results among different LVLMs (lower values demonstrate fewer hallucinations).
 With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination.
-However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination [5, 6, 10] (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning.
+However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination (Li et al., 2023e; Wang et al., 2023b; Chen et al., 2024c) (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning (Gou et al., 2025a).
 As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image.
 Therefore, a unified definition and taxonomy is necessary to integrate different kinds of LVLM hallucination.
-2) Hallucination discrimination: existing works [5, 6, 8, 11] rely on self-discrimination (e.g., Yes/No questions) to testify LVLMs’ awareness of objects and relations in the given image.
-When prompted with an instruction “Is there a {content} in the image?”, hallucination is determinied only when LVLMs answer “Yes.” for a non-existing content.
-However, these benchmarks require transforming general vision-language tasks into "yes/no" formats, limiting their applicability.
-Additionally, different LVLMs may have different capability in answering such“yes/no” questions since they are pre-trained on different data, which may make the evaluation results biased [12].
-So, here arises our question: can we perform a unified and unbiased evaluation for various kinds of hallucination?
-To this end, this paper first proposes a unified framework to measure both object and relation hallucinations (§3).
-Specifically, it first extracts knowledge graphs as triplets from LVLMs’ responses, then uses external judges to compare these triplets with scene graphs of input images for hallucination evaluation.
-In this way, our method can provide fine-grained evaluation results on the response to any vison-language tasks.
-Based on the proposed framework, we then construct Tri-HE, a novel Triplet-level Hallucination Evaluation benchmark to evaluate object and relation hallucination simultaneously (§4).
-As reported in §5 and Figure 1(b), relation hallucination is indeed a common nightmare even for the powerful GPT-4V model [13], and is even more severe than object hallucination.
-By comparing the performance of different LVLMs, we further recognize several findings that could potentially decrease LVLM hallucination (§5.2).
-In addition, our triplet-level hallucination judge empowered by GPT-4 reveals remarkable alignment with human judgments (Table 3).
-This motivates us to integrate triplet description into LVLMs and propose a simple yet effective training-free hallucination mitigation method (§5.4).
-With this method, the LLaVA-1.5 model [14] exceeds all the open-sourced counterparts, achieving comparable performance with the powerful GPT-4V, as shown in Figure 1(b) and Table 5.
-Our main contributions contain three parts:
+2) Hallucination discrimination: To evaluate how severe LVLMs hallucinate objects and their relationships within given images, prior works generally use either self-discrimination methods (e.g., Yes/No questions) (Li et al., 2023e; Wang et al., 2023b; Guan et al., 2024; Wu et al., 2024b) or template-driven discrimination approaches (e.g., “What is the relation with A and B?”) such as Reefknot (Zheng et al., 2024).
+However, such methods inherently constrain LVLMs to generate short answers like “Yes/No” or “A has {} relation to B”.
+Given that LVLMs have varying capabilities to produce brief responses due to differences in pre-training datasets, this could introduce biased evaluation results (Chen et al., 2023b; Liu et al., 2024).
+For instance, Li et al.
+(2023e) have shown that InstructBLIP (Dai et al., 2023) tends to produce shorter outputs compared to other LVLMs, thus inflating its performance in answering the above type of questions and leads to hallucination evaluation bias.
+Moreover, these benchmarks require transforming general vision-language tasks into specific formats like "Yes/No", limiting their applicability.
+Therefore, we raise the following research question: Can we develop a unified and unbiased evaluation framework capable of evaluating various types of hallucinations in LVLM responses across diverse tasks?
+To this end, we first propose a unified framework to simultaneously measure object and relation hallucinations in LVLM responses (§3).
+Specifically, our framework extracts knowledge graphs represented as triplets from LVLM-generated responses and then employs external evaluators to compare these triplets against the corresponding scene graphs from the input images.
+Consequently, our method facilitates hallucination evaluation for responses across diverse vision-language tasks, independent of the specific question formats.
+Leveraging this unified framework, we further introduce the Tri-HE, a novel benchmark for Triplet-level Hallucination Evaluation, designed explicitly to assess both object and relation hallucinations (§4).
+Our experimental findings presented in §5 and Figure 1(b) confirm that relation hallucination poses a significant challenge for both closed-source and open-source LVLMs, often surpassing object hallucination in severity.
+By systematically comparing LVLMs’ performance, we identify key insights that could potentially reduce hallucination rates (§5.2).
+Furthermore, our proposed triplet-level hallucination judge, powered by LLMs, demonstrates impressive alignment with human judgments (Table 3).
+Motivated by these observations, we incorporate explicit triplet descriptions into LVLM prompts and introduce a straightforward yet effective training-free method to mitigate hallucinations (§5.4).
+Our primary contributions are summarized in the following three perspectives:
   1. 1.
-    We introduce a unified framework to measure both the object and relation hallucination of LVLMs simultaneously.
-Specifically, we propose to conduct a triplet-level hallucination evaluation, which can provide more accurate and fine-grained evaluation results.
+    We propose a unified framework capable of jointly evaluating object and relation hallucinations in LVLM responses across diverse vision-language tasks.
+In particular, our triplet-level evaluation offers a finer-grained, more accurate assessment compared to existing methods.
   2. 2.
-    Based on the framework, we propose Tri-HE, a novel triplet-level fine-grained hallucination evaluation benchmark for LVLMs.
+    Building upon this framework, we introduce Tri-HE, a novel triplet-level fine-grained hallucination evaluation benchmark tailored specifically for LVLMs.
   3. 3.
-    We propose a simple yet effective training-free hallucination mitigation method that exceeds all open-sourced counterparts, performing on par with GPT-4V on Tri-HE.
-(a) Triplet-level LVLM hallucination evaluation pipeline.
-(b) LVLM hallucination comparison.
-Figure 1: Overview of Tri-HE evaluation pipeline.
-(a) With the provision of images, scene graphs, and questions, knowledge graphs (i.e., triplets) are extracted from LVLM responses, which are then judged by GPT-4.
-(b) The radar plot showcases the evaluation results among different LVLMs (lower values demonstrate fewer hallucinations).
+    We propose a simple yet highly effective training-free hallucination mitigation approach that surpasses the open-source LVLM competitors.
 2 Related Work
@@
 LLMs are empowered to understand and reason about both images and text by aligning representations from visual encoders to pre-trained language models, followed by visual instruction tuning.
-LLaVA [14, 2] proposes to use a simple projection layer to integrate the visual representations into textual encoders, which is further enhanced in Shikra [15] by incorporating referential dialogue tasks.
-Instead, BLIP [16] proposes the Q-Former architecture to extract useful information from the visual representations, which is also used by MiniGPT-4 [3] and InstructBLIP [1].
-MoCLE [17] further introduces the Mixture-of-Experts architecture into LVLMs to deal with the data conflict during instruction tuning.
-Although powerful, existing works primarily focus on improving the helpfulness, without a thorough analysis on the reliability of LVLMs.
+LLaVA (Liu et al., 2023a; b) proposes to use a simple projection layer to integrate the visual representations into textual encoders, which is further enhanced in Shikra (Chen et al., 2023d) by incorporating referential dialogue tasks.
+Instead, BLIP (Li et al., 2023a) proposes the Q-Former architecture to extract useful information from the visual representations, which is also used by MiniGPT-4 (Zhu et al., 2023) and InstructBLIP (Dai et al., 2023).
+InternLM (Dong et al., 2024) aligns with more diverse instruction data with the conditional online reinforcement learning from human feedback (RLHF) strategy, while MoCLE (Gou et al., 2023) further introduces the Mixture-of-Experts architecture into LVLMs to deal with the data conflict during instruction tuning.
+Although powerful, existing works primarily focus on improving the helpfulness and robustness Gou et al.
+(2025b), without a thorough analysis of the reliability of LVLMs.
 2.2 Hallucination Evaluation in LVLMs
-With the prevalence of LVLMs, a growing number of studies have been conducted on their hallucination issues [9, 18, 19, 20, 21, 6, 8, 22].
-Previous hallucination evaluation works can be categorized into two groups: 1) solely evaluating object hallucinations or do not distinguish different hallucinations [23, 24, 6, 10], which neglects other hallucination types like relation hallucination and is thus not comprehensive.
-The other type of works use “yes/no” questions to evaluate LVLM’s relation/object hallucinations [5, 25, 8, 11].
+With the prevalence of LVLMs, a growing number of studies have been conducted on their hallucination issues (Chen et al., 2024b; d; Han et al., 2024; Huang et al., 2023; Li et al., 2023b; Wang et al., 2023b; Guan et al., 2024; Yue et al., 2024).
+Previous hallucination evaluation works can be categorized into two groups: 1) solely evaluating object hallucinations or do not distinguish different hallucinations (Zhao et al., 2023; Li et al., 2023c; Wang et al., 2023b; Chen et al., 2024c), which neglects other hallucination types like relation hallucination and is thus not comprehensive.
+The other type of works use “yes/no” questions to evaluate LVLM’s relation/object hallucinations (Li et al., 2023e; Wang et al., 2023a; Guan et al., 2024; Wu et al., 2024b).
 However, these benchmarks require transforming general vision-language tasks into “yes/no” formats, limiting their applicability.
@@
 To remedy this research gap, our paper proposes a triplet-level evaluation framework that can provide fine-grained object and relation hallucinations for responses to any vision-language tasks, with an evaluation benchmark Tri-HE that incorporates questions requiring more complicated commonsense reasoning.
-It is worth noting that a concurrent benchmark, Reefknot [26], also evaluates relation hallucinations at the triplet level.
-However, compared to Tri-HE, Reefknot has several limitations: 1) It fully relies on existing datasets, which may lead to data contamination, whereas Tri-HE consists of questions synthesized by GPT-4V.
-2) Reefknot evaluates only a limited number of relation types, while Tri-HE covers a broader range, enabling more comprehensive evaluation results as shown in §5.
-3) Reefknot employs only one entailment-based hallucination discriminator, whereas Tri-HE offers more accurate hallucination detection using GPT-4, which can simultaneously identify both object and relation hallucinations.
-3 Tri-HE Framework Formulation
-Inspired by the relation extraction [27] tasks in NLP, in this section, we propose a unified framework to evaluate both object and relation hallucinations via the object-relation triplets (i.e., $(\text{Object}_{1},\text{Relation},$ $\text{Object}_{2})$).
+It is noteworthy that a concurrent benchmark, Reefknot (Zheng et al., 2024), similarly assesses relation hallucinations at the triplet level.
+However, Reefknot exhibits several limitations compared to Tri-HE.
+First, Reefknot constructs VQA questions based on a simple template, “What is the relation between A and B?”, restricting both the variety of vision-language tasks that can be evaluated and the length of LVLM-generated responses, potentially introducing evaluation biases, similarly with “yes/no” questions for evaluating object hallucination.
+In contrast, our framework is flexible enough to be applied to various vision-language tasks.
+Moreover, since the questions in Tri-HE are generated by GPT-4V, it can cover a wider range of relation types compared to template-based questions, thus providing more comprehensive evaluation results.
+Second, Reefknot relies solely on a single entailment-based hallucination discriminator, whereas Tri-HE leverages powerful LLM-based discriminators capable of accurately and simultaneously identifying both object and relation hallucinations, leading to more comprehensive hallucination evaluation results.
+3 Unified Hallucination Evaluation Framework Formulation
+Inspired by the relation extraction (Xiaoyan et al., 2023) task, in this section, we propose a unified framework to evaluate both object and relation hallucinations via the object-relation triplets (i.e., $(\text{Object}_{1},\text{Relation},$ $\text{Object}_{2})$).
 Here the objects and relations can either be a word or a phrase with attributes.
@@
 3.1 Definitions
-As shown in Figure 1(a), given an input image $I$, a question $Q$ related to $I$ with its ground truth answer $A$, and the answer $A_{\theta}$ predicted by a LVLM $A_{\theta}(\cdot|Q,I)$ parameterized with $\theta$, we first define,
+As illustrated in Figure 1(a), we formulate our framework with the standard VQA setting (although they can be generalized to evaluate hallucinations in any vision-language tasks given available scene graph annotations, as discussed in §3.4).
+Specifically, considering an input image $I$, a corresponding question $Q$ associated with image $I$, its ground truth answer $A$, and the answer $A_{\theta}$ predicted by an LVLM $A_{\theta}(\cdot|Q,I)$ parameterized by $\theta$, we can first have the following definitions:
   • $G=(V,E)$ as the scene graph of $I$, where $V$ and $E$ refer to all the objects existing in $I$ and all the possible relations among existing objects, respectively.
-  • $G^{\prime}=(V^{\prime}\subseteq V,E^{\prime}\subseteq E)$ as the knowledge graph that includes all the required objects and relations to answer $Q$.
+  • $G^{\prime}=(V^{\prime}\subseteq V,E^{\prime}\subseteq E)$ as the knowledge graph that includes all the required objects and relations to answer $Q$ precisely.
   • $G_{\theta}=(V_{\theta},E_{\theta})$ as the knowledge graph extracted from $A_{\theta}$, where $V_{\theta}$ and $E_{\theta}$ include all the objects and all the possible relations among objects mentioned in $A_{\theta}$.
 Note that here all graphs can be converted to a set of triplets (i.e., $G=\{(v_{1},e,v_{2})\}$, where $v_{1},v_{2}\in V$ and $e\in E$).
-A common nightmare in previous LVLM hallucination literature lies in the ambiguous discrimination between prediction hallucinations and errors [28].
+A common nightmare in previous LVLM hallucination literature lies in the ambiguous discrimination between prediction hallucinations and errors (Ji et al., 2023).
 To obtain unbiased hallucination evaluation results, we separate them depending on whether or not the wrongly generated objects or relations exist in the given image $I$.
@@
   • Object hallucination: if $v_{1}\notin V$ or $v_{2}\notin V$, suggesting $A_{\theta}$ includes an object not within $I$.
-For example, the triplet (location, suggests, popular spot for socializing) in Figure 1(a) encounter an object hallucination since the object “ popular spot for socializing” cannot be obtained from $V$.
+For example, the triplet (location, suggests, popular spot for socializing) in Figure 1(a) encounters an object hallucination since the object “ popular spot for socializing” cannot be obtained from $V$.
   • Relation hallucination: if $v_{1},v_{2}\in V$ yet $e\notin E$, suggesting that $A_{\theta}$ correctly recognizes two related objects from $I$ but pair them with a non-existing relation.
@@
 3.2 Evaluation Metrics
-With the above definition in hand, given the knowledge graph $G_{\theta}$ extracted from a model response $A_{\theta}$, we calculate hallucination rates of $A_{\theta}$ as the proportion of hallucinated triplets in $G_{\theta}$.
-Most previous works (e.g., POPE [5]) directly evaluate the hallucination rate at the object-level with respect to the total number of predicted objects, yet make their results not comparable among LVLMs, since different LVLMs might refer to different numbers of objects in their responses.
-To address this issue, we instead opt to calculate the hallucination rate in the question- and image-level.
+With the above definition in hand, given the knowledge graph $G_{\theta}$ extracted from a model response $A_{\theta}$, we calculate the hallucination rates of $A_{\theta}$ as the proportion of hallucinated triplets in $G_{\theta}$.
+Most previous works (e.g., POPE (Li et al., 2023e)) directly evaluate the hallucination rate at the object-level with respect to the total number of predicted objects, yet make their results not comparable among LVLMs, since different LVLMs might refer to different numbers of objects in their responses.
+To address this issue, we instead opt to calculate the hallucination rate at the question- and image-level.
 Specifically, we calculate two types of hallucination rates, including the question-level hallucination rate ($\text{Hallu}_{\text{Q}}$) and image-level hallucination rate ($\text{Hallu}_{\text{I}}$), as defined in the following,
@@
 3.3 Evaluation Pipeline
-With the definitions and evaluation metrics provided in §3.1 and §3.2, the remaining problems contain two parts: 1) how to extract the knowledge graph $G_{\theta}$ from LVLM responses $A_{\theta}$, and 2) how to judge a triplet in $G_{\theta}$ is hallucinated or not.
+With the definitions and evaluation metrics provided in §3.1 and §3.2, the remaining problems contain two folds: 1) how to extract the knowledge graph $G_{\theta}$ from LVLM responses $A_{\theta}$, and 2) how to judge whether a triplet in $G_{\theta}$ is hallucinated or not.
 The overview of our pipeline is illustrated in Figure 1(a).
@@
 Given an LVLM response $A_{\theta}$ with the corresponding question $Q$ and image $I$, we extract the knowledge graph $G_{\theta}$ from $A_{\theta}$ via prompting GPT-4.
-Check our prompt for knowledge graph extraction in Figure 6.
-Afterward, we propose two different strategies to judge whether a triplet $(v_{1},e,v_{2})\in G_{\theta}$ includes hallucination based on the ground truth answer $A$ and the image scene graph $G$, as described in the following.
-Figure 2: Visualization of data samples in Tri-HE. Each image is associated with a scene graph and question-answer pairs with the reasoning triplet annotations.
+Check our prompt for knowledge graph extraction in Appendix §A.1.
+Afterwards, we propose two different strategies to judge whether a triplet $(v_{1},e,v_{2})\in G_{\theta}$ includes hallucination based on the ground truth answer $A$ and the image scene graph $G$, as described in the following.
 NLI Judge.
-The first strategy is implemented with a natural language inference (NLI) [29] model  [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
+The first strategy is implemented with a natural language inference (NLI) (Reimers & Gurevych, 2019) model  [^1] https://huggingface.co/sentence-transformers/all-mpnet-base-v2.
 Specifically, given an extracted triplet, we first calculate its cosine similarity scores with all triplets in the image scene graph $G$ and only retain those ground truth (GT) triplets with similarity scores greater than 0.5 to refine the information that will be used for the NLI model.
@@
 If the NLI score between the extracted triplet and ground truth triplets is lower than 0.6, suggesting the extracted triplet cannot be induced based on GT triplets, and therefore, resulting in a hallucination.
-# Images # Questions # Objects # Relations # Questions / Image # Triplets / Scene Graph
+To determine the threshold, we randomly selected question instances from 10 images and reviewed the set of filtered triplets that were returned.
+The similarity score threshold was adjusted to 0.5 for the most reasonable returned triplets.
+These triplets later concatenate together as the ground truth required for generating NLI judgments.
+In determining if a generated triplet was hallucinated, we further review the NLI judgment results in different thresholds, ultimately deciding on a threshold of 0.6.
+LLM Judge.
+Another evaluation strategy is to leverage prompting of a powerful LLM, a widely-adopted practice in recent works for assessing LLM outputs (Zheng et al., 2023).
+In this work, we primarily utilize GPT-4 in LLM judge to determine whether a given extracted triplet $(v_{1},e,v_{2})\in G_{\theta}$ can be directly obtained or inferred from the image scene graph $G$.
+Note that:
+  1. 1.
+    We do not employ GPT-4V in LLM judge, as Li et al. (2024) have reported that the text-only GPT-4 judge is more consistent with human preferences than GPT-4V judge.
+  2. 2.
+    Frontier open-source models, such as LLaMA-3.3, can similarly deliver reliable and cost-efficient hallucination evaluation results (see detailed analysis in Table 3), which might be superior alternatives when the GPT-4 judge is unavailable.
+Additionally, if a triplet $(v_{1},e,v_{2})$ is judged as hallucinated, we further prompt the LLM judge to clarify whether the hallucination pertains specifically to the relation $e$ or the objects $v_{1},v_{2}$.
+Refer to Appendix §A.2 for the prompt of LLM judge in our experiments  [^2] For both knowledge graph extraction and LLM judge, we utilize the “gpt-4-1106-preview” model via OpenAI’s API with default inference parameters..
+3.4 Generalizability of our Framework
+Although we formulate our unified triplet-level hallucination evaluation framework in the sections mentioned above primarily based on the VQA tasks, it is capable of evaluating hallucinations in LVLM responses for any vision-language task (since knowledge graph extraction is suitable for all natural-language-based responses), provided that the corresponding scene graphs for the test images are available or extracted by pre-trained expert models.
+This underscores the task-agnostic design of our proposed framework and highlights its strong generalization capability.
+We leave the detailed exploration for other vision-language tasks for future work.
+4 Tri-HE Construction
+Following the formulation in §3, in this section, we provide a detailed discussion on how to construct our benchmark Tri-HE for a unified triplet-level evaluation of both hallucinations in LVLMs.
+Image Collection.
+The construction of Tri-HE begins with images from the GQA dataset (Hudson & Manning, 2019), as the scene graph annotations provided by GQA naturally fit our triplet-level hallucination evaluation formulation.
+Nevertheless, some scene graphs in GQA contain incomplete object relationships, omitting information necessary for accurate question answering.
+To mitigate this issue, we initially filter the GQA images, retaining only those whose scene graphs contain at least five object relations (edges between nodes).
+Subsequently, we select 300 images from the filtered images according to the following criteria:
+  1. 1.
+    Each image must contain more than two related objects.
+  2. 2.
+    Each image must be sufficiently clear to discern all visual details.
+This procedure ensures a set of high-quality images suitable for subsequent dataset construction.
+VQA Question Generation
+Next, since the VQA questions in the GQA dataset have already been extensively used during the pre-training of current LVLMs, we instead employ GPT-4V  [^3] We use the “gpt-4-vision-preview” model here, the same as in §5.2.
+to generate novel question-answer pairs for each image to avoid data contamination.
+To effectively examine both object and relation hallucinations in LVLM responses, we aim to generate questions that necessitate commonsense reasoning grounded on the provided images.
+Specifically, for every image, GPT-4V is prompted to generate 10 questions along with their answers  [^4] Check §A.3 for the prompt used here., each requiring image-based commonsense reasoning to be answered.
+Furthermore, we ask GPT-4V to produce relation triplets describing the reasoning processes of answering the questions.
+These additional triplets can subsequently be used to enrich the original image scene graphs.
+VQA Question Verification
+Following the initial generation of VQA questions, three annotators manually examine all generated questions, answers, and triplets based on the following criteria:
+  1. 1.
+    Each question must be valid and answerable using commonsense reasoning based on the input image.
+  2. 2.
+    Each answer must appropriately address the question using commonsense reasoning.
+  3. 3.
+    Each triplet must accurately describe the corresponding answer and must only contain objects visible within the image.
+Questions or answers failing to meet these conditions are discarded, while invalid triplets are excluded from the respective scene graphs.
+To validate the annotation consistency, the annotators jointly annotate 100 common question-answer pairs, achieving a Krippendorff’s alpha (Krippendorff, 2011) of 0.62, demonstrating substantial inter-annotator agreement and guaranteeing the consistency of Tri-HE annotations.
+Statistics.
+The overall statistics for Tri-HE are summarized in Table 1.
+As described in Figure 2, each image in Tri-HE is linked to a scene graph and a set of question-answer pairs that require reasoning, accompanied by ground truth triplet annotations.
+Note that since the quality of each question in Tri-HE is manually verified, expanding its size requires significant resources and poses challenges.
+Nonetheless, the number of images and questions in Tri-HE is comparable to existing LVLM hallucination evaluation benchmarks such as Zhao et al.
+(2023) and Guan et al.
+(2024).
+Furthermore, as demonstrated in Table 3, Tri-HE has been sufficient to provide reliable hallucination evaluation results, largely thanks to the high-quality annotation procedure described above.
+#Images #Questions #Objects #Relations #Questions/Image #Triplets/SG
 300 1226 1723 618 4.09 19.10
-Table 1: Statistics of Tri-HE.
-GPT-4 Judge.
-The other strategy is via prompting the GPT-4 model, which is a common practice for recent LLM evaluation literature [30].
-We prompt GPT-4 to justify whether a previously extracted triplet $(v_{1},e,v_{2})\in G_{\theta}$ can be directly obtained or inferred based on the image scene graph $G$.
-Note that we do not adopt the GPT-4V judge, since [31] have observed that the text-only GPT-4 judge better aligns with the human preference than GPT-4V.
-Moreover, if $(v_{1},e,v_{2})$ is recognized as a hallucinated triplet, we further require it to tell whether the hallucination relates to the relation $e$ or objects $v_{1},v_{2}$.
-Check our prompt for GPT-4 judge in Figure 8  [^2] For both knowledge graph extraction and GPT-4 judge, we use the “gpt-4-1106-preview” model via OpenAI’s API calls with its default inference parameters..
-With triplet-level hallucination judgments in hand, we can adopt Eq.
-(1) and Eq.
-(2) to calculate the corresponding hallucination rates of $A_{\theta}$, which can be further separated into object hallucination and relation hallucination rates if GPT-4 judge is adopted.
-Check more detailed results in Table 2.
-4 Tri-HE Construction
-Following the formulation in §3, in this section, we provide a detailed discussion on how to construct our Tri-HE benchmark for a unified triplet-level evaluation of both hallucinations in LVLMs.
-Dataset.
-We build Tri-HE upon the GQA dataset [32], which requires LVLMs to conduct commonsense reasoning based on a given image.
-The provision of scene graph annotations aligns with our formulation of hallucination evaluation at the triplet-level.
-Data collection.
-However, several scene graphs in GQA are only equipped with parts of the relations among objects appearing in the image, making necessary information required for question answering absent.
-To address that, we first employ a filtering process to identify images with adequate distinct relations in the scene graph annotation.
-Specifically, we start by selecting 300 images with more than 5 relations from the corresponding GQA scene graph annotations.
-To enlarge the question-answer pairs for each GQA image, we utilize GPT-4V  [^3] We use the “gpt-4-vision-preview” model here, as in §5.2.
-to generate new questions requiring commonsense reasoning based on the input image.
-By leveraging the prompt provided in Figure 7, we prompt GPT-4V to generate ten questions for each image along with their corresponding answers and relation triplets that illustrate the reasoning process.
-Afterward, we manually verify the correctness of generated questions, answers, and triplets, while only questions indeed requiring reasoning are maintained.
-Moreover, the verified triplets are added to the original scene graphs collected in GQA.
-Method GPT-4 Judge NLI Judge
+Table 1: Statistics of Tri-HE. “SG” refers to Scene Graph.
+Figure 2: Visualization of data samples in Tri-HE. Each image is associated with a scene graph and question-answer pairs with the reasoning triplet annotations.
+5 Evaluation Results
+5.1 Evaluated LVLMs
+We selected six open-source LVLMs for evaluation, including the LLaVA series (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023), Shikra (Chen et al., 2023d), and InternLM-XComposer2 (abbrev., InternLM-X2) (Cai et al., 2024).
+For all evaluated LVLMs, we selected the 7B variants to ensure fair comparison.
+Additionally, we test the recent popular Llama-3.2-Vision-Instruct (abbrev., LLaMA-3.2) (MetaAI, 2024a) and use its smallest version (11B).
+The prompt templates and inference configurations used for LVLMs are detailed in Appendix §A.4 and §B.
+All experiments are conducted on two Nvidia A100 GPUs.
+5.2 Main Result
+LVLM comparison.
+Table 2 compares hallucination rates of different LVLMs on our Tri-HE benchmark.
+As can be seen, all the evaluated LVLMs suffer from generating hallucinations with at least 38% hallucination rates.
+Among these LVLMs, InternLM-X2 (Cai et al., 2024) obtains the best overall performances, suggesting that its strategy to train with both text-image and textual-only instruction data simultaneously helps better align its visual encoder and LLM, and thus, reduces its hallucination rates.
+Moreover, compared to LLaVA (Liu et al., 2023b), Shikra (Chen et al., 2023d) has consistently lower hallucination rates, which is built upon LLaVA’s structure with extra grounding capability introduced, indicating that introducing extra grounding could help LVLMs reduce hallucination.
+Additionally, LLaMA-3.2 achieves the lowest relation hallucination rates, suggesting that a strong textual backbone can help mitigate relation hallucination.
+However, it exhibits a weaker ability to accurately identify objects, impacting its object and overall hallucination rates.
+Since LLaMA-3.2 does not outperform other LVLMs with even more parameters, we do not adopt it in the remaining experiments for parameter consistency.
+Relation hallucination is more severe.
+Except for MiniGPT-4 and LLaMA-3.2, all the LVLMs generate more relation hallucinations than object hallucinations.
+A possible explanation is that existing LVLMs lack reasoning abilities, which makes them easily confused and mess up the relations among objects.
+This further suggests that focusing on object hallucination (Li et al., 2023e) is not enough for a throughout analysis of the LVLM reliability, and a unified and comprehensive study like our triplet-level evaluation is necessary.
+Evaluation pipeline.
+In addition, we observe that LLM Judge can provide clearer and more reasonable discrimination between models compared to the NLI judge.
+We provide a more comprehensive investigation into the differences between these two judges later in §5.3.
+Besides, the evaluation results under both $\text{Hallu}_{\text{I}}$ and $\text{Hallu}_{\text{Q}}$ metrics demonstrate similar trends, proving the robustness of our proposed triplet-level hallucination evaluation setting under different evaluation granularities.
+Method LLM Judge NLI Judge
 Overall Object Relation Overall
@@
 LLaMA-3.2 40.16 38.95 22.30 21.08 17.86 17.87 48.46 45.64
-InternLM2 38.83 37.54 18.25 17.50 20.58 20.04 54.41 52.08
-Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM2 is short for InternLM-XComposer2 [33]
-Statistics.
-The overall statistics for Tri-HE are summarized in Table 1.
-As described in Figure 2, each image in Tri-HE is linked to a scene graph and a set of question-answer pairs that require reasoning, accompanied by ground truth triplet annotations.
-Note that since the quality of each question in Tri-HE is manually verified, expanding its size requires significant resources and poses challenges.
-Nonetheless, the number of images and questions in Tri-HE is comparable to existing LVLM hallucination evaluation benchmarks such as [23] and [8].
-Furthermore, as demonstrated in §5, Tri-HE is able to produce reliable hallucination evaluation results.
-5 Results on the Tri-HE Benchmark
-5.1 Evaluated LVLMs
-We selected six open-source LVLMs for evaluation, including the LLaVA series [14, 2], MiniGPT-4 [3], InstructBLIP [1], Shikra [15], and InternLM-XComposer2 (abbrev., InternLM2) [33].
-For all evaluated LVLMs, we selected the 7B variants to ensure fair comparison.
-Additionally, we test the recent popular Llama-3.2-Vision-Instruct model (abbrev., LLaMA-3.2) [34] and used its smallest version (11B).
-The official prompt templates and inference configurations were used for all LVLMs.
-All experiments are conducted on two Nvidia A100 GPUs.
-5.2 Main Result
-LVLM comparison.
-Table 2 compares hallucination rates of different LVLMs on our Tri-HE benchmark.
-As can be seen, all the evaluated LVLMs suffer from generating hallucinations with at least 38% hallucination rates.
-Among these LVLMs, InternLM2 [33] obtains the best overall performances, suggesting that its strategy to train with both text-image and textual-only instruction data simultaneously helps better align its visual encoder and LLM, and thus, reduces its hallucination rates.
-Moreover, compared to LLaVA [2], Shikra [15] has consistently lower hallucination rates, which is built upon LLaVA’s structure with extra grounding capability introduced, indicating that introducing extra grounding could help LVLMs reduce hallucination.
-Additionally, LLaMA-3.2 achieves the lowest relation hallucination rates, suggesting that a strong textual backbone can help mitigate relation hallucination.
-However, it exhibits a weaker ability to accurately identify objects, impacting its object and overall hallucination rates.
-Since LLaMA-3.2 does not outperform other LVLMs with even more parameters, we do not adopt it in the remaining.
-Relation hallucination is more severe.
-Except for MiniGPT-4 and LLaMA-3.2, all the LVLMs generate more relation hallucinations than object hallucinations.
-A possible explanation is that existing LVLMs lack reasoning abilities, which makes them easily confused and mess up the relations among objects.
-This further suggests that focusing on object hallucination [5] is not enough for a throughout analysis of the LVLM reliability and a unified and comprehensive study like our proposed triplet-level evaluation is necessary.
-Evaluation pipeline.
-In addition, we observe that GPT-4 judge can provide clearer and more reasonable discrimination between models compared to NLI judge.
-We provide a more comprehensive investigation into the differences between these two judges later in §5.3.
-Besides, the evaluation results under both $\text{Hallu}_{\text{I}}$ and $\text{Hallu}_{\text{Q}}$ metrics demonstrate the same trend, proving the robustness of our proposed triplet-level hallucination evaluation setting under different evaluation granularities.
-Method LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM2 GPT-4V
+InternLM-X2 38.83 37.54 18.25 17.50 20.58 20.04 54.41 52.08
+Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM-X2 is short for InternLM-XComposer2 (Cai et al., 2024)
+Method LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM-X2 GPT-4V
 NLI Judge (Sentence) 0.2182 0.0970 0.3609 0.2596 0.2684 0.2524 0.2787
 NLI Judge (Triplet) 0.2951 0.2838 0.2264 0.4259 0.2829 0.2647 0.4190
-GPT-4 Judge (Sentence) 0.6631 0.5409 0.3669 0.5532 0.5821 0.5998 0.5548
-GPT-4 Judge (Triplet) 0.8115 0.6320 0.4283 0.6235 0.6939 0.7169 0.7292
-Table 3: Pearson correlation scores among automatic hallucination judgments and human judgments. The best results under each column are boldfaced.
-LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM2
+LLM Judge (Llama-3.3 Sentence) 0.4705 0.4842 0.3617 0.2520 0.4941 0.4366 0.4969
+LLM Judge (Llama-3.3 Triplet) 0.5138 0.5262 0.4150 0.4798 0.5311 0.5323 0.5519
+LLM Judge (GPT-4 Sentence) 0.6631 0.5409 0.3669 0.5532 0.5821 0.5998 0.5548
+LLM Judge (GPT-4 Triplet) 0.8115 0.6320 0.4283 0.6235 0.6939 0.7169 0.7292
+Table 3: Pearson correlation scores among automatic hallucination judgments and human judgments.
+The best results under each column are boldfaced.
+The LLM Judges utilized are specified in the brackets.
+Evaluate Closed-sourced LVLMs.
+In addition to evaluating open-sourced LVLMs, we further investigate the performance of closed-sourced LVLMs.
+Due to limited experimental resources, we specifically evaluate the GPT-4V (OpenAI, 2023) model on a subset of 25 randomly selected images from Tri-HE.
+Specifically, we prompt GPT-4V to obtain responses to all questions related to these selected images and compute the associated hallucination rates following the steps described in Table 2.
+For comparison, we also include results from open-sourced LVLMs evaluated on the same set of 25 images.
+As illustrated in Figure 1(b), GPT-4V clearly demonstrates superior performance, surpassing all open-sourced LVLMs.
+Although GPT-4V exhibits slightly higher object hallucination rates compared to InternLM-X2—likely because it tends to associate additional objects not present in the image—it achieves notably lower relation hallucination rates due to its stronger reasoning capabilities, resulting in lower overall hallucination rates.
+5.3 Analysis
+Investigating automatic hallucination judgments with human judgments.
+Here, we further illustrate the effectiveness of the triplet-level evaluation setting by studying its correlation with human judgments.
+To conduct fine-grained hallucination analysis, previous works (Jing et al., 2023; Min et al., 2023) split a model response into sub-sentences first, on which their hallucination measurements are conducted.
+We regard this method as a baseline for comparison.
+Specifically, we sample a subset of 20 images from Tri-HE and invite human annotators to score five-point-scale hallucination rates of the responses of all the LVLMs in §5.1 (check Appendix §C for the detailed annotation guidelines).
+The human annotators achieve a Krippendorff’s alpha score of 0.66, indicating a high inter-agreement.
+Results are shown in Table 3.
+We find that triplet-level hallucination rates have higher correlations with human judgments with both NLI and LLM Judges, indicating that identifying hallucination on triplets can lead to a more accurate, human-preferred evaluation for model responses.
+Moreover, we notice that the LLM Judges achieves a higher correlation to human judgments compared to the NLI counterpart, revealing LLMs’ superior abilities to find hallucinations, which is also consistent with our observation in §5.2.
+LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM-X2
 Original $\text{Hallu}_{\text{I}}\downarrow$ 22.46 22.03 25.25 24.50 23.65 20.58
@@
 $\text{Hallu}_{\text{Q}}\downarrow$ 18.73 18.06 22.68 19.82 19.34 16.10
-Table 4: Relation hallucination rates for the top 20% frequent object pairs of different LVLMs under the GPT-4 judge. Original refers to the results in Table 2.
-Performance of GPT-4V.
-Except for the open-sourced LVLMs, we also explore the performance of state-of-the-art commercial LVLMs [13, 35].
-Due to the experiment budget, here we only evaluate the GPT-4V [13] model with 25 randomly selected samples in Tri-HE.
-Specifically, we directly prompting GPT-4V to gather its answers and calculate the corresponding hallucination rates, following the same setting in Table 2.
-We also report the performances of open-sourced LVLMs on the 25 selected images.
-As can be seen in Figure 1(b), GPT-4V demonstrates its superior capability and outperforms all open-sourced LVLMs.
-Although its object hallucination rates are slightly higher than InternLM2 since it tends to associate many objects that are not in the image, GPT-4V has significantly lower reasoning hallucination rates than other LVLMs, indicating its strong reasoning ability, and therefore, reports superior overall hallucination rates.
-5.3 Analysis
-Investigating automatic hallucination judgments with human judgments.
-In §3, we propose to measure hallucination on triplet-level and design two automatic hallucination judges.
-Here, we further illustrate the effectiveness of the triplet-level evaluation setting by studying its correlation with human judgments.
-To conduct fine-grained hallucination analysis, previous works [36, 37] split a model response into sub-sentences first, on which their hallucination measurements are conducted.
-We regard this method as a baseline for comparison.
-Specifically, we sample a subset of 20 images from Tri-HE and invite human annotators to score five-point-scale hallucination rates of the responses of all the LVLMs in §5.1.
-Check Table 9 for the detailed annotation guidelines.
-The human annotators achieve a Krippendorff’s alpha score [38] of 0.66, indicating a high inter-agreement.
-Results are shown in Table 3.
-We find that triplet-level hallucination rates have higher correlations with human judgments with both NLI and GPT-4 judges, indicating that identifying hallucination on triplets can lead to a more accurate, human-preferred evaluation for model responses.
-Moreover, we notice that the GPT-4 judge achieves a higher correlation to human judgments compared to the NLI counterpart, revealing GPT-4’s superior ability to find hallucinations, which is also consistent with our observation in §5.2.
-Figure 3: An illustration demonstrating hallucination mitigation.
-The three prompting strategies (w/o Mitigation, General Description, and Triplet Description) are listed from left to right.
-Hallucinated content is highlighted in Red and repeating contents are marked with italic and underline.
+Table 4: Relation hallucination rates for the top 20% frequent object pairs of different LVLMs under the LLM Judge. Original refers to the results in Table 2.
+Applying Different LLMs in LLM Judge.
+While GPT-4 allows the LLM Judge to produce reliable hallucination evaluations, the associated API expenses could become large when evaluating a large number of examples.
+To mitigate potential cost constraints, we also examine whether alternative open-source LLMs can serve effectively in LLM Judge.
+Specifically, we replace GPT-4 with LLaMA-3.3-70B-Instruct (abbrev., Llama-3.3) (MetaAI, 2024b) and re-evaluate all examples listed in Table 3.
+As shown, similar to GPT-4, Llama-3.3 consistently achieves higher correlation scores at the triplet-level than at the sentence-level.
+Furthermore, its Pearson correlation scores with human evaluations, while significantly outperforming those obtained using NLI Judge, remain comparable to GPT-4’s results for certain LVLMs.
+These findings suggest that open-source LLMs can serve as viable alternatives to GPT-4 in LLM Judge, providing reliable evaluation results under tight budget constraints, thereby further validating the robustness of our proposed LLM Judge.
 Investigating relation hallucination with object information.
 As concluded from §5.2, existing LVLMs tend to generate both object and relation hallucinations in their replies, while the relation hallucination rates are even higher.
-Since different LVLMs have pairs of objects $(v_{1},v_{2})$ that they are familiar with (e.g., high-frequency object pairs in the instruction data they are fine-tuned on) and might generate correct relations on these objects easily, we suppose that the relation hallucination problem might mostly be located in less-frequent object pairs.
-To verify this assumption, we extract all object pairs for each LVLM from their respective $G_{\theta}$ generated from responses on Tri-HE, and rank these pairs based on their frequency  [^4] Details of this process can be seen in Appendix D..
+Since different LVLMs have pairs of objects $(v_{1},v_{2})$ that they are familiar with (e.g., high-frequency object pairs in the instruction data they are fine-tuned on) and might generate the correct relations on these objects easily, we suppose that the relation hallucination problem might mostly be located in less-frequent object pairs.
+To verify this assumption, we extract all object pairs for each LVLM from their respective $G_{\theta}$ generated from responses on Tri-HE, and rank these pairs based on their frequency.
 Then, we calculate each LVLM’s relation hallucination rates on their most frequent object pairs.
-As in Table 4, all the LVLMs have significantly lower relation hallucination rates on frequent object pairs they are familiar with, suggesting that they know the possible relations among objects and understand how to choose a relation appropriately.
-Check more analyses regarding the impact of input length on hallucination rates in Appendix A.
-Mitigation GPT-4 Judge NLI Judge
-$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
-w/o Mitigation 45.86 47.44 55.93 54.94
-General Description 46.50 49.19 54.59 53.03
-Triplet Description 44.14 42.96 51.19 47.12
-Table 5: Hallucination mitigation results on MiniGPT-4. The best results under each column are boldfaced.
-5.4 Hallucination Mitigation
-Regarding the reasons of LVLM hallucination, previous studies [36, 39] have identified modality misalignment as a key factor.
-Inspired by ECSO [40], we propose a training-free LVLM hallucination mitigation method via self-alignment [12, 40, 41].
-Specifically, we disable the direct visual access of LVLMs to alleviate hallucinations resulting from modality misalignment and fully leverage instruction-following capabilities of LVLMs’ (text-only) LLM backbone for hallucinations mitigation [24].
-Method.
-We start by prompting LVLMs to describe the image on awareness of the given question (i.e., General Description).
-Subsequently, we prevent LVLMs from having direct access to the image to downgrade the text-image LVLMs to text-only LLMs, which are then encouraged to respond only based on the previously generated image description but without access to the image.
-As discussed in §5.3, better hallucination evaluation is observed at the triplet level.
-Thus, we further prompt LVLMs to focus more on identifying objects and inter-object relations (i.e., Triplet Description) as triplets when describing the image.
-We use the subset in Figure 1(b) with MiniGPT-4, with the prompts and an example in Figure 3.
-Results.
-As shown in Table 5, our results demonstrate the effectiveness of hallucination mitigation through our proposed prompting strategies.
-Both general and triplet descriptions achieve a significant hallucination rate reduction while prompting LVLMs to focus on objects and inter-object relations achieves the lowest hallucination rates.
-Additionally, we conduct ablation studies and compare our method with existing baselines in Appendix B.
-6 Conclusion
-Starting from a unified definition of hallucinations, we propose a novel triplet-level LVLM hallucination evaluation framework for both object and relation hallucinations.
-Then we introduce Tri-HE, a novel triplet-level LVLM hallucination evaluation benchmark, with which, we conduct a throughout analysis of the discrepancy among object and relation hallucinations.
-Finally, we propose a simple yet effective training-free hallucination mitigation method, which integrates our findings regarding objects and inter-object relations.
-Limitations
-Due to the experiment budget, in Figure 1(b), we evaluate GPT-4V on a randomly selected subset of 25 images; in Table 3, we conduct human annotation on another randomly selected subset of 20 images.
-Although we do not interfere with the data selection process, there may still exist a few biases in the corresponding results.
-Also, we do not have the budget to evaluate more LVLMs on Tri-HE.
-We will try to provide more comprehensive results in future works once we get more experimental budgets.
-Appendix
-Appendix A More Analyses
+To obtain object pairs and their rankings from LVLM responses, suppose that for the targeted LVLM, we have its responses to all questions in Tri-HE, i.e., $G_{\theta}$, we first extract all the object pairs $(v_{1},v_{2})$ from $G_{\theta}$.
+Then, for each object, we replace it with the name of its synset using WordNet to reduce the total types of objects.
+Afterward, we could calculate the frequency of each object pair and rank them based on their frequency.
+This ranking will then be used to calculate the first 20% frequent object pairs in Table 4.
+As in Table 4, all the LVLMs have significantly lower relation hallucination rates on frequent object pairs they are familiar with, suggesting that they know the possible relations among objects and understand how to choose a relation appropriately for frequently occurring objects.
 Investigating hallucination rates with response length.
-Previous studies on LVLM hallucination evaluation suggest that the length of model responses may influence the extent of hallucination [5, 39], as some LVLMs tend to produce shorter, safer outputs.
+Previous studies on LVLM hallucination evaluation suggest that the length of model responses may influence the extent of hallucination (Li et al., 2023e; Zhou et al., 2023), as some LVLMs tend to produce shorter, safer outputs.
 However, directly instructing an LVLM to generate a response of a specific length is challenging.
@@
 Figure 3: Trends of the hallucination rates of the image-level (left) and question-level (right) evaluations for different LVLMs with respect to the number of tokens in the model responses.
-Appendix B More on Hallucination Mitigation
+5.4 Hallucination Mitigation
+After demonstrating that LVLMs exhibit significant hallucination problems, we further explore potential approaches to reduce both object and relation hallucinations.
+Prior works (Jing et al., 2023; Zhou et al., 2023; Li et al., 2023c; Gou et al., 2024) have suggested that modality misalignment might be a primary cause behind LVLM hallucinations.
+Motivated by this claim, we propose a training-free method to mitigate hallucinations by improving modality alignment within LVLMs.
+Figure 4: An illustration demonstrating hallucination mitigation.
+The three prompting strategies (w/o Mitigation, General Description, and Triplet Description) are listed from left to right.
+Hallucinated content is highlighted in Red and repeating contents are marked with italic and underline.
+Method.
+Specifically, we propose a two-step strategy.
+Given an image and its corresponding question, we first prompt the evaluated LVLM to generate a description of the image guided by the given question (General Description in Figure 4).
+Next, we prompt the same LVLM (in a new version without the image memory) using this generated description to answer the question.
+Through this approach, we effectively leverage the strong instruction-following capability intrinsic to the LVLM’s LLM backbone, instead of requiring the LVLM to simultaneously comprehend the image and answer the question, thereby reducing hallucinations caused by modality misalignment.
+Moreover, as indicated in §5.3, triplet-level evaluation is more effective than sentence-level evaluation in assessing hallucinations.
+Hence, we further explicitly guide LVLMs to concentrate more on identifying objects and their interrelations via triplets when describing images (Triplet Description in Figure 4).
+We evaluate MiniGPT-4 and LLaVA-1.5 combined with our proposed mitigation approaches using the subset previously employed in Figure 1(b).
+The corresponding prompts along with an example illustration are shown in Figure 4.
+Mitigation LLM Judge NLI Judge
+$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
+MiniGPT-4 w/o Mitigation 45.86 47.44 55.93 54.94
+General Description 46.50 49.19 54.59 53.03
+Triplet Description 44.14 42.96 51.19 47.12
+LLaVA-1.5 w/o Mitigation 30.72 30.17 53.84 52.06
+General Description 28.70 29.80 51.40 49.80
+Triplet Description 28.39 32.68 48.97 48.40
+Table 5: Hallucination mitigation results. The best results under each column are boldfaced.
+Results.
+As demonstrated in Table 5, both LVLMs exhibit reduced hallucination rates after applying our two-stage mitigation method, indicating that improved modality alignment effectively alleviates hallucinations.
+In addition, explicitly prompting LVLMs to emphasize objects and their relationships consistently yields the lowest hallucination rates across most cases, further reinforcing the findings presented in Table 3.
 Ablation Study.
-We conducted an ablation study on a general description (i.e., “describe the image”).
-As shown in Table 6, the combination of both triplet description prompting and disability of visual input contributed to the effectiveness of our hallucination mitigation approach.
-Mitigation GPT-4 Judge NLI Judge
+We also perform an ablation study on the best-performing “Triplet Description” variant of our mitigation approach to gain deeper insights into the role of each module within our proposed method.
+Specifically, we compare the Triplet Description (i.e., Triplet+Eyes-Close) results obtained by MiniGPT-4 with two alternative setups:
+  1. 1.
+    Eyes-Close: This setting is equivalent to General Description.
+Image access is disabled (i.e., eyes-close) while prompting LVLM to answer the question.
+It is designed to assess the impact of employing triplet-level descriptions.
+  2. 2.
+    Triplet: This setting is similar to Triplet Description but allows image accessibility.
+It incorporates both the original image and the generated triplet-level description simultaneously as inputs.
+It is designed to examine the effects of modality alignment.
+The experimental results are presented in Table 6.
+As shown, the combined use of triplet-level description and restriction of visual input access leads to the lowest hallucination rates.
+These findings further validate the design choices made in our mitigation method.
+Mitigation LLM Judge NLI Judge
 $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
 w/o Mitigation 45.86 47.44 55.93 54.94
-Ours (eyes-close) 46.50 49.19 54.59 53.03
-Ours (triplet) 45.65 45.16 59.35 55.57
-Ours (triplet+ eyes-close) 44.14 42.96 51.19 47.12
-Table 6: Ablation study on MiniGPT-4 [3]. The best results under each column are boldfaced.
-Mitigation on LLaVA-1.5.
-To provide a comprehensive evaluation, we conducted additional experiments using another LVLM (i.e., LLAVA-1.5) to mitigate hallucinations.
-The results presented in Table 7 indicate that both the general description and triplet description methods show an improved reduction in hallucinations, further highlighting the effectiveness of our mitigation approach.
-Mitigation GPT-4 Judge NLI Judge
-$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
-w/o Mitigation 30.72 30.17 53.84 52.06
-General Description 28.70 29.80 51.40 49.80
-Triplet Description 28.39 32.68 48.97 48.40
-Table 7: Further experiment conduct on LLaVA-1.5 [14]. The best results under each column are boldfaced.
-Baseline Comparison.
-Current methods for mitigating hallucinations typically involve retraining, integrating external detection models, and devising decoding strategies.
-Compared to existing works, our approach is a plug-and-play method that neither requires costly retraining nor relies on external models.
-To make a more fair comparison, we conducted experiments with LogicCheckGPT [42], a training-free approach that addresses hallucinations through prompting interactions with the help of GPT-3.5.
-Under the cost consideration, an evaluation was conducted only with the NLI judge.
-The results indicate that our method outperforms LogicCheckGPT, highlighting its effectiveness in mitigating hallucinations as shown in Table 8.
+Eyes-Close 46.50 49.19 54.59 53.03
+Triplet 45.65 45.16 59.35 55.57
+Triplet+ Eyes-Close 44.14 42.96 51.19 47.12
+Table 6: Ablation study on hallucination mitigation. The best results under each column are boldfaced.
 Model Mitigation $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
@@
 Ours (triplet+ eyes-close) 51.19 47.12
-Table 8: Hallucination mitigation results on LLaVA-1.5 [14] and MiniGPT-4 [3] with baseline comparison. The best results under each column are boldfaced.
-Qualitative comparison
-with LogicCheckGPT [42] is provided in Figure 5.
-Figure 5: More illustrations demonstrating hallucination mitigation with comparison to LogicCheckGPT. Hallucinated content is highlighted in Red.
-Appendix C NLI Threshold Selection
-We randomly selected question instances from 10 images and reviewed the set of filtered triplets that were returned.
-The similarity score threshold was adjusted to 0.5 for the most reasonable returned triplets.
-These triplets later concatenate together as the ground truth required for generating NLI judgments.
-In determining if a generated triplet was hallucinated, we further review the NLI judgment results in different thresholds, ultimately deciding on a threshold of 0.6.
-Appendix D Object Pairs Extraction and Ranking
-In this section, we detailedly describe how we obtain object pairs and their rankings from LVLM responses.
-Suppose we have all an LVLM’s responses to all questions in Tri-HE, i.e., $G_{\theta}$, we first extract all the object pairs $(v_{1},v_{2})$ from $G_{\theta}$.
-Then for each object, we replace it by the name of its synset using WordNet to reduce the total types of objects.
-Afterward, we could calculate the frequency of each object pair and rank them based on their frequency.
-This ranking will then be used to calculate the first 20% frequent object pairs in Table 4.
-Appendix E More on Prompts
-E.1 Prompt for triplets extraction with GPT-4
+Table 7: Hallucination mitigation results on LLaVA-1.5 Liu et al. (2023a) and MiniGPT-4 Zhu et al. (2023) with baseline comparison. The best results under each column are boldfaced.
+Baseline Comparison.
+Current methods for mitigating hallucinations typically involve retraining, integrating external detection models, and devising decoding strategies.
+Compared to existing works, our approach is a plug-and-play method that neither requires costly retraining nor relies on external models.
+To make a more fair comparison, we conducted experiments with LogicCheckGPT Wu et al.
+(2024a), a training-free approach that addresses hallucinations through prompting interactions with the help of GPT-3.5.
+Under the cost consideration, an evaluation was conducted only with the NLI judge.
+The results indicate that our method outperforms LogicCheckGPT, highlighting its effectiveness in mitigating hallucinations as shown in Table 7.
+Qualitative comparison with LogicCheckGPT (Wu et al., 2024a) is provided in Figure 8.
+6 Conclusion
+Starting from a unified definition of hallucinations, we propose a novel triplet-level LVLM hallucination evaluation framework for both object and relation hallucinations.
+Then we introduce Tri-HE, a novel triplet-level LVLM hallucination evaluation benchmark, with which, we conduct a throughout analysis of the discrepancy among object and relation hallucinations.
+Finally, we propose a simple yet effective training-free hallucination mitigation method, which integrates our findings regarding objects and inter-object relations.
+Acknowledgments
+This work has been made possible by a Research Impact Fund project (RIF R6003-21) and a General Research Fund project (GRF 16203224) funded by the Research Grants Council (RGC) of the Hong Kong Government.
+Appendix A Prompts
+A.1 Prompt for triplets extraction with GPT-4
 The prompt for extracting triplets in the answer generated by LVLMs is illustrated in Figure 5.
@@
 Figure 5: Prompt for triplets extraction with GPT-4.
-E.2 Prompt for question generation with GPT-4V
-The prompt for generating questions, answers, and corresponding triplets with GPT-4V is shown in Figure 7.
-Generate ten questions about the given image that require an inferential answer, which is not directly observable from the image.
-The answer to each question can be explained by one or more (object, relation, object) triplets that appear in the scene graph of the given image.
-Note that the triplets should consist of objects and relations that are visible in the given image.
-Output the results in the format of: Generated Questions: Answers: Explanations:
-Figure 7: Prompt for question generation with GPT-4V.
-E.3 Prompt for GPT-4 Judge
-The prompt for our proposed GPT-4 judge method is illustrated in Figure 8.
+A.2 Prompt for LLM Judge
+The prompt for our proposed LLM judge method is illustrated in Figure 6.
 Given a list of reference triplets ("object1", "relation", "object2") extracted from the scene graph of an image, along with a list of objects observed in this image, your task is: Task 1.
@@
 Reference: <REFERENCE> List of Objects: <LIST_OF_OBJECTS> Claim: <CLAIM>
-Figure 8: Prompt for the GPT-4 Judge method.
-E.4 Prompt for LVLMs in Evaluation
-The prompt for generating responses from LVLMs for evaluation is the combination of a question and the corresponding image.
-Appendix F Human Annotation Guideline
+Figure 6: Prompt for the LLM Judge method.
+A.3 Prompt for question generation with GPT-4V
+The prompt for generating questions, answers, and corresponding triplets with GPT-4V is shown in Figure 7.
+Generate ten questions about the given image that require an inferential answer, which is not directly observable from the image.
+The answer to each question can be explained by one or more (object, relation, object) triplets that appear in the scene graph of the given image.
+Note that the triplets should consist of objects and relations that are visible in the given image.
+Output the results in the format of: Generated Questions: Answers: Explanations:
+Figure 7: Prompt for question generation with GPT-4V.
+A.4 Prompts for Evaluating LVLMs
+When evaluating LVLMs on Tri-HE, the prompt we use is the question itself. Questions are fed into LVLMs along with the corresponding images.
+Appendix B Configurations for LVLM Evaluation
+For LVLM evaluations, we directly use the default configuration settings provided in their publicly available code repositories.
+For instance, the configurations utilized for evaluating LLaVA models are accessible at https://github.com/haotian-liu/LLaVA.
+Appendix C Human Annotation Guideline
 The detailed guidelines of our human evaluation tasks are shown in Table 8. Noting that two types of inferences in the model responses are regarded as hallucinations during human annotation:
@@
 Table 8: Detailed human evaluation instructions.
-Appendix G More Discussion
+Figure 8: More illustrations demonstrating hallucination mitigation with comparison to LogicCheckGPT. Hallucinated content is highlighted in Red.
+Appendix D More Discussion
 Future works.
-Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models [43, 44, 45, 46, 47, 48] is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder [49, 50, 51, 52] and visual tools (e.g., object detectors [53, 54]) to better extract visual information for LLM reasoning.
+Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models (Chen et al., 2023c; Gao et al., 2023; 2024a; 2024b; Li et al., 2023d; Liu et al., 2023c; Wang et al., 2024) is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder (Chen et al., 2021; 2023a; Liu et al., 2022; Zhili et al., 2023) and visual tools (e.g., object detectors (Han et al., 2021; Li et al., 2022)) to better extract visual information for LLM reasoning.

### v4

@@
 With the prevalence of LVLMs, enormous works have started to explore the evaluation and analysis of LVLM hallucination.
-However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination (Li et al., 2023e; Wang et al., 2023b; Chen et al., 2024c) (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting commonsense reasoning (Gou et al., 2025a).
+However, two problems are observed: 1) Hallucination category: most existing works focus on object-related hallucination (Li et al., 2023e; Wang et al., 2023b; Chen et al., 2024c) (i.e., LVLMs describing an object not existing in the given image) while ignoring the possibility that even when two objects are successfully recognized, LVLMs might still mess up with their relationships when conducting reasoning over these objects(Gou et al., 2025a; Yu et al., 2025).
 As illustrated in the example in Figure 1(a), LLaVA successfully recognizes the “people” and the train station “area”, yet predicts their relation to be “walking around” that cannot be directly obtained from the given image.
@@
 However, such methods inherently constrain LVLMs to generate short answers like “Yes/No” or “A has {} relation to B”.
-Given that LVLMs have varying capabilities to produce brief responses due to differences in pre-training datasets, this could introduce biased evaluation results (Chen et al., 2023b; Liu et al., 2024).
+Given that LVLMs have varying capabilities to produce brief responses due to differences in pre-training datasets, this could introduce biased evaluation results (Chen et al., 2023b; Liu et al., 2024b).
 For instance, Li et al.
@@
 InternLM-X2 38.83 37.54 18.25 17.50 20.58 20.04 54.41 52.08
-Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE. The best results under each column are boldfaced. InternLM-X2 is short for InternLM-XComposer2 (Cai et al., 2024)
+Table 2: Comparison on hallucination rates among different LVLMs on Tri-HE.
+The best results under each column are boldfaced.
+InternLM-X2 is short for InternLM-XComposer2 (Cai et al., 2024).
+Check Appendix D for evaluation results on more LVLMs.
 Method LLaVA LLaVA-1.5 MiniGPT-4 InstructBLIP Shikra InternLM-X2 GPT-4V
@@
 Ours (triplet+ eyes-close) 51.19 47.12
-Table 7: Hallucination mitigation results on LLaVA-1.5 Liu et al. (2023a) and MiniGPT-4 Zhu et al. (2023) with baseline comparison. The best results under each column are boldfaced.
+Table 7: Hallucination mitigation results on LLaVA-1.5 Liu et al.
+(2023a) and MiniGPT-4 Zhu et al.
+(2023) with baseline comparison.
+The best results under each column are boldfaced.
+Check more baseline comparison results in Appendix D.
 Baseline Comparison.
@@
 Figure 8: More illustrations demonstrating hallucination mitigation with comparison to LogicCheckGPT. Hallucinated content is highlighted in Red.
-Appendix D More Discussion
-Future works.
-Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models (Chen et al., 2023c; Gao et al., 2023; 2024a; 2024b; Li et al., 2023d; Liu et al., 2023c; Wang et al., 2024) is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder (Chen et al., 2021; 2023a; Liu et al., 2022; Zhili et al., 2023) and visual tools (e.g., object detectors (Han et al., 2021; Li et al., 2022)) to better extract visual information for LLM reasoning.
+Appendix D Additional Results
+In this section, we present additional evaluation results to supplement those in Table 2 and Table 5, enabling a more comprehensive evaluation.
+Due to the deprecation of gpt-4-1106-preview—previously employed in LLM Judge (GPT-4 Triplet)—by OpenAI’s API, and our limited budget for querying proprietary models, we adopt LLM Judge (Llama-3.3 Triplet) for the following experiments.
+Its effectiveness has been demonstrated in Table 3.
+D.1 Evaluating more recent LVLMs on Tri-HE
+To ensure a more comprehensive evaluation, we further evaluate InternVL2_5-8B Chen et al.
+(2024e) and Qwen2.5-VL-7B-Instruct Team (2025) on Tri-HE.
+Since we use the LLM Judge (LLaMA-3.3 Triplet) in this setting—which differs from the judge used in Table 2—we also report the results of InternLM-X2 and LLaMA-3.2 under the same judge for comparison.
+The results are presented in Table 9.
+As shown, all recent models still exhibit a certain degree of both object and relation hallucinations, demonstrating the effectiveness of Tri-HE in identifying hallucination issues in LVLMs.
+Notably, InternVL2_5-8B achieves the lowest hallucination rates across all metrics, suggesting that its superior pre-training data quality and the use of Random JPEG Compression as a data augmentation strategy are effective in mitigating hallucinations in model responses.
+Method LLM Judge
+Overall Object Relation
+$\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$ $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
+LLaMA-3.2 38.98 37.37 23.25 21.13 13.51 13.30
+Qwen2.5-VL-7B-Instruct 38.19 35.92 26.16 24.09 11.83 11.38
+InternLM-X2 31.75 29.93 19.29 18.09 12.46 11.84
+InternVL2_5-8B 29.12 26.99 17.16 14.95 12.22 11.13
+Table 9: Comparison on hallucination rates among recent LVLMs on Tri-HE with LLM Judge (Llama-3.3 Triplet).
+D.2 More Hallucination Mitigation Methods
+In this section, we include additional recent training-free methods for hallucination reduction to enable a more comprehensive comparison.
+Specifically, we compare our mitigation framework with a decoding-based approach, VCD (Leng et al., 2024).
+As shown in Table 10, although VCD is designed to address object hallucination, it exhibits more severe hallucination issues when evaluated on our reasoning-grounded question set compared to our proposed mitigation method.
+In addition, reinforcement learning (RL) approaches have received growing attention for addressing hallucination.
+To enable a more comprehensive evaluation, we assess a RL-trained method, OPA-DPO (Yang et al., 2025), using our evaluation framework.
+The results in Table 10 demonstrate its effectiveness.
+It is worth noting that, since OPA-DPO requires additional training, its performance advantage over our method is expected.
+Nonetheless, it can potentially be integrated with our proposed approach as a foundation for further RL fine-tuning.
+Model Mitigation Type $\text{Hallu}_{\text{I}}\downarrow$ $\text{Hallu}_{\text{Q}}\downarrow$
+LLaVA-1.5 w/o Mitigation - 21.73 21.07
+VCD training-free 28.89 30.64
+Ours training-free 19.53 20.98
+OPA-DPO RL-trained 14.44 16.82
+Table 10: Additional baseline comparisons on hallucination mitigation with Llama3.3 Judge. The best results under each column are boldfaced.
+Appendix E Future works
+Currently, the proposed triplet-level evaluation is primarily deployed on LVLMs, whose extension to diffusion models (Chen et al., 2023c; Gao et al., 2023; 2024a; 2024b; Li et al., 2023d; Liu et al., 2024a; Wang et al., 2024) is feasible, while for the hallucination mitigation proposed in §5.4 can be further enhanced by utilizing stronger vision encoder (Chen et al., 2021; 2023a; Liu et al., 2022; Zhili et al., 2023) and visual tools (e.g., object detectors (Han et al., 2021; Li et al., 2022)) to better extract visual information for LLM reasoning.
+Furthermore, additional types of hallucination can be incorporated into our triplet-level evaluation framework, such as temporal changes of an object over time, represented in the form of (old, change, new) triplets.
+We plan to explore these extensions in future work.